### v2

Abstract
We study a full implementation problem where a socially desirable outcome depends on a state of the world which is unknown to the designer but (commonly) known to a set of agents. The designer may ask the agents to present hard evidence which they privately draw from a distribution depending on the state. We identify a necessary and sufficient condition for implementation in (mixed-strategy) Bayesian Nash equilibria called No Perfect Deceptions (NPD). When agents also face uncertainty about the state and are informationally small (McLean and Postlewaite (2002)), a generalization of the NPD condition (GNPD) is sufficient for implementation and necessary when the information size is zero. Our implementing mechanisms accommodate the case with two or more agents, invoke no integer/modulo games, and impose transfers that in equilibrium vanish with the agents’ information size.
1 Introduction
Consider a government allocating the rights to an oil drilling project amongst some firms. Depending on the scale of the deposits, the government wants to allocate the drilling rights to one of the firms. The firms receive private but correlated signals about the deposits through their prospecting operations, and may also be able to provide some (stochastically available or uncertain) evidence regarding these signals. The government wishes to estimate the scale of the deposits based on the signals received by the firms, but firms do not necessarily have the incentive to report their signals truthfully since each firm wants the drilling rights regardless of the scale of the deposits.
Alternately, consider a personal injury trial, in which a judge must determine appropriate compensation from the defendant to the plaintiff. The judge seeks to match compensation to actual damages but lacks direct knowledge of them. Both parties know the true damage, but have opposing interests: the plaintiff wants higher compensation, the defendant lower. Evidence comes from investigation and is therefore uncertain, though both parties may anticipate what evidence might emerge.
These examples share several key features. First, the designer lacks knowledge of a payoff-relevant state, $s$. Second, agents have preferences that are independent of the state but potentially misaligned with those of the designer. Third, there exists hard evidence in the form of messages that agents can send, which are only available in certain states of the world and thus refute other states. However, the availability of this evidence is uncertain – it is stochastic – so agents are unsure about the evidence others may possess. Finally, monetary incentives are feasible.
The two examples differ in one crucial aspect: while trial participants have common knowledge of damages, competing firms face limited uncertainty about each other’s extraction costs (their “state types”). [^1] Such uncertainty will be modeled using one of two approaches - nonexclusive information (Postlewaite and Schmeidler (1986)) and information smallness (McLean and Postlewaite (2004)). Such settings are widespread, making it essential to develop robust mechanisms for credible information elicitation.
In studying full implementation when hard evidence is available, the literature has so far focused mainly on the case of deterministic evidence (Kartik and Tercieux (2012), Ben-Porath and Lipman (2012), Banerjee, Chen, and Sun (2024b)) and complete information about the decision relevant state. [^2] Peralta (2019) studies Bayesian implementation in a setting with evidence, but focuses on direct mechanisms. We allow for indirect mechanisms which do not feature integer or modulo games. While this yields positive results under weak conditions, evidence is often the result of research studies which are fundamentally uncertain in nature, so the assumption that agents have complete information in evidence endowments is often unrealistic, motivating the study of uncertain evidence.
When studying implementation problems with uncertain evidence, several possibilities emerge with respect to the choice of information structure regarding the agents’ knowledge of the decision relevant state of the world. The examples above show that the agents often share significant information, even when the designer lacks it. In the trial, parties commonly know the damages. Similarly, oil drilling firms likely have substantial knowledge about competitors’ capabilities. In the literature, McLean and Postlewaite (2002) discuss information smallness, which states that aggregating all but one firm’s information largely reveals the remaining firm’s type [^3] Loosely speaking, information smallness refers to the notion that the influence of any one agent on the posterior belief about the state (in this case the profile of state types) is small., and Postlewaite and Schmeidler (1986) propose that such aggregation yields perfect revelation of the remaining state type (a condition termed Nonexclusive Information, or NEI).
Thus, in this paper, we study how credible decisions may be made when the designer cannot rely on preference variation, and also cannot expect the agents to have deterministic evidence to separate states. Rather, evidence is uncertain, and the information structure may involve complete information, nonexclusive information, or information smallness in the decision relevant state of the world. We hold that mechanisms must fully implement, i.e. achieve the socially desired outcome in each equilibrium.
We obtain the following results. First, under common knowledge, a condition we term No Perfect Deceptions (NPD) characterizes implementable SCFs without transfers. An SCF satisfies NPD if it prescribes identical outcomes in two states $s$ and $s^{\prime}$ whenever there is a profile of strategies (a deception) in $s$ which mimic the equilibrium strategies in $s^{\prime}$ while there is no article of evidence at $s$ which refutes $s^{\prime}$. Surprisingly, a stochastic version of the measurability condition in Ben-Porath and Lipman (2012) does not suffice. [^4] Under pure strategies, the weaker No Pure Perfect Deceptions (NPPD) condition applies, though still stronger than stochastic measurability from Ben-Porath and Lipman (2012). This clarifies that in this setting, mixed strategy implementation is strictly harder than pure strategy implementation. The reader is referred to the online appendix for more details. Second, with informationally small agents, a generalization, Generalized NPD (GNPD) suffices for implementation with zero expected transfers in large economies. Third, under NEI, GNPD characterizes all implementable SCFs without transfers. These conditions are readily satisfiable – small perturbations to the evidence structure make such deceptions refutable by evidence.
We establish the sufficiency of these conditions by constructing implementing mechanisms which use evidence variation in three distinct ways. First, in presenting evidence, an agent directly refutes all the states where she may not possess such evidence with positive probability. Second, if an agent cannot present the evidence for a state claim (perhaps because some evidence is not available with enough probability), the distribution induced on her evidence report varies from the expected distribution under truth-telling - a difference which we will show can be used to flag incorrect reporting using a system of bets. Third, if such a bet is placed to blow the whistle on an incorrect report profile, we incentivize the agents to present all their evidence (using a monetary transfer which rewards them for each article they present) and use a system of scoring rules which asks agents to predict the (truthful) evidence presentation of the others. With sufficient evidence variation, this scoring rule extracts the truth in the state dimension, and this can be used to realign the report profile toward the truth by a system of crosschecks.
The mechanisms we construct make several important methodological contributions. First, unlike most Bayesian implementation results which rely on problematic integer or modulo games, our mechanisms achieve implementation through incentive alignment. [^5] For a description of the issues with integer and modulo games, we refer the reader to Jackson (1992). We eschew the canonical approach of first obtaining consistent reports using an integer game and then arguing that consistent reports may only obtain on the truth. Instead, when agents deviate from truth-telling, we detect this through properties of the message profile and use transfer rules to realign incentives toward truthful reporting, which is closer to the intended spirit of mechanism design. Additionally, while most Bayesian mechanisms require three or more agents and restrict attention to pure strategies, our implementation works with just two agents and allows mixed strategies.
Second, we apply techniques from partial implementation literature to full implementation. We adapt betting systems similar to those in Crémer and McLean (1988) and proper scoring rules as in Johnson, Pratt, and Zeckhauser (1990) and McLean and Postlewaite (2015). While using scoring rules for exact implementation is challenging due to their on-path transfers, we overcome this by combining them with crosschecks that cancel out transfers in equilibrium. Under truthful evidence presentation, scoring rules elicit true state types from agents predicting others’ evidence, while crosschecks ensure that we avoid equilibrium transfers. [^6] With complete or nonexclusive information, crosschecks yield no transfers. Under the weaker information smallness condition, transfers occur with small probability but are zero in expectation, while maintaining incentive compatibility as in McLean and Postlewaite (2002). Further, we find bets using a separating hyperplane argument in the spirit of Crémer and McLean (1988), but use them as a whistleblowing device for agents to signal that a deception has been played, as opposed to their use in Crémer and McLean (1988) for efficient implementation with correlation. For technical details, we refer the reader to Section 4.4.
Finally, a key innovation addresses implementation under uncertain state types, where higher-order beliefs are crucial since social outcomes in states $s$ and $s^{\prime}$ must be identical under GNPD when agents cannot distinguish between them at any belief order. We develop a technique to embed higher-order beliefs directly into the message space, enabling evidence-based incentivization even when bets are expected only at high belief orders. [^7] This technique is useful in settings where the designer wishes to incentivize certain actions on the part of the agents based on their higher order beliefs.
These techniques – the equilibrium cancellation of scoring rules and the elicitation of higher-order beliefs into the message space, arguably add important tools to the field of mechanism design, especially when the canonical approach involving integer games is not used. For instance, the latter finds use in Banerjee, Chen, and Sun (2024a), where we study implementation of SCFs in correlated equilibria. These techniques address a common challenge: incentivizing agent behavior using monetary transfers based on low-probability events while avoiding transfers on the equilibrium path.
We extend our analysis to scenarios where agents face uncertainty about the state, and evidence draws may be correlated. In this context, we demonstrate that implementation in rationalizable strategies is achievable with small transfers alongside when at least one agent can differentiate between states through their belief hierarchies (which we term “higher-order measurability” or HOM), enabling implementation of non-constant social choice functions even without preference variation.
The rest of the paper is organized as follows. In Section 2, we formally describe an illustrating example to provide context for the analysis in later sections. Section 3 defines the setup formally and introduces some properties of the evidence structure. Section 4 deals with the Bayes-Nash implementation result. Section 5 discusses implementation with informational smallness and nonexclusive information. Section 6 details rationalizable implementation in the more general setup. Finally, Section 7 reviews and compares our results with the literature.
2 Illustrating Example: Drilling for Oil
Consider a government allocating drilling rights for an oil tract. The deposit contains either a low ($L$), medium ($M$), or high ($H$) quantity of oil. Three firms are competing for the rights: firm $A$ (a small wildcatter), firm $B$ (an established local firm), and firm $C$ (a large corporation). Each firm conducts its own preliminary assessment, receiving a signal that could be low ($l$), medium ($m$), or high ($h$). The states of the world are given by triads like $lll$, $lml$ etc. These signals are correlated, but the correlation structure varies depending on the information environment.
Each firm’s assessment provides a signal about the deposit size and also generates evidence of varying conclusiveness. While a firm receiving a low signal can only present inconclusive test results, a firm receiving a medium or high signal might obtain evidence definitive enough to rule out a small deposit. The government’s objective is to allocate the drilling rights efficiently: to firm A if the deposit is small (due to their cost advantages), to firm B if medium-sized (given their local expertise), and to firm C if large (reflecting their large-scale capabilities). The government plans to use majority rule on the signals reported to a mechanism to infer the deposit size, defaulting to treating it as small if no majority emerges. Each firm wants the rights regardless of the state of the world.
The evidence structure reflects the varying conclusiveness of assessments. When a firm’s tests indicate a high deposit (signal $h$), there is a 60% chance these tests are definitive enough to rule out the possibility of a small deposit. In this case, they can present evidence $E=\{m,h\}$, indicating their tests show the deposit must be either medium or large. With 40% chance, their tests are less conclusive, yielding evidence $E=\{l,m,h\}$, consistent with any deposit size.
Similarly, when a firm receives a medium signal ($m$), there is a 40% chance their assessment can definitively rule out a small deposit (yielding evidence $E=\{m,h\}$), and a 60% chance their tests are inconclusive (yielding evidence $E=\{l,m,h\}$). When a firm’s tests indicate a small deposit (signal $l$), their evidence is always inconclusive ($E=\{l,m,h\}$), reflecting the technical difficulty of definitively ruling out larger deposits through preliminary testing. Formally, we can represent this evidence structure as shown below.
Signal Received Evidence Distribution
$h$ $0.6\times\{\{m,h\},\{l,m,h\}\}+0.4\times\{\{l,m,h\}\}$
$m$ $0.4\times\{\{m,h\},\{l,m,h\}\}+0.6\times\{\{l,m,h\}\}$
$l$$\{\{l,m,h\}\}$
Note that the articles of evidence in the above example have been associated with subsets of the signal space. In general, we could have abstract articles of evidence, such as $e_{1}$ and $e_{2}~{}$instead of $\{l,m,h\}$, and $\{m,h\}$ respectively. These “names” are obtained by associating an article of evidence with the states in which they appear with positive probability. This form of nomenclature is used in Ben-Porath and Lipman (2012), albeit in a deterministic setting. We use this nomenclature to simplify the exposition (especially within examples) in several places in the paper. Note that in this setting, types contain the signal received and the evidence endowment, e.g. $(m,\{l,m,h\})$.
From the above intuition of evidence, it is clear that no agent can eliminate state $mmm$ at state $hhh$. This poses a major challenge in achieving implementation. To see this, suppose the mechanism were direct, requiring only a state and an evidentiary claim. Further, consider the following strategies for any firm A of type $(h,\{\{m,h\},\{l,m,h\}\})$: Report $\{\{m,h\},\{l,m,h\}\}$ with probability $\frac{2}{3}$ and report $\{\{l,m,h\}\}$ with probability $\frac{1}{3}$ in the evidence dimension; In the state dimension, always claim that the received signal is $m$. With this strategy, the agents would induce the same distribution on their reports as if the true state were $mmm$, and they were reporting truthfully. Such a strategy is later dubbed a perfect deception. If perfect deceptions exist, it is not possible to separate states in any mechanism (direct or indirect) which relies on Bayesian Nash Implementation. This result is formally proved in Section 4.1 and followed by an intuitive description. Note that there is also a perfect deception from state $mmm$ to state $lll$, but $lll$ is refutable at $mmm$ with positive probability (by the article of evidence $\{m,h\}$), so that states $mmm$ and $lll$ are separable by a mechanism.
[…]
Turning to the informational structure of the received signals, there may be several possibilities. Following McLean and Postlewaite (2002) or McLean and Postlewaite (2004), if agents are informationally small, and if a majority of the firms receive the same signal ($l$ for instance), then with high probability, the remaining firm also receives the same signal $l$. Nonexclusive information (Postlewaite and Schmeidler (1986), Vohra (1999)) holds that if two firms receive the same signal, the third firm must receive the same signal. Finally, the notion of complete information in this setting implies that the scale of the deposits and the signals are perfectly correlated, so that if one firm gets a signal ($m$ for instance), then the remaining firms must also get the same signal. We will study full implementation for each of these information structures alongside uncertain evidence. Finally, note that while in this example it is likely that firms have evidence only for their own signals, in other settings it may be the case that agents may have evidence about the state types (signals) of other agents as well, and the model we study allows for this.
3 Model and Preliminaries
3.1 Setup
There is a set of agents =$\mathcal{I=}\left\{1,...,I\right\},$ $I\geq 2$, a set of outcomes $A$ and a finite set of states $S.$ Agents have bounded utility over the outcomes given by $\bar{u}_{i}:A\times S\rightarrow\mathbb{R}$. Their preferences are quasilinear, so that $u_{i}(a,\bar{\tau}_{i},s)=\bar{u}_{i}(a,s)+\bar{\tau}_{i}$, where $\bar{\tau}_{i}$ is the transfer of money to agent $i$. The bound of $\bar{u}_{i}$ can be normalized so that it is strictly less than $1$, that is, an agent can be persuaded to accept any outcome if the alternative were any other outcome with a penalty of $1$ dollar.
Each agent $i$ is endowed with a collection of articles of evidence $E_{i}$ which can vary from state to state. A profile of collections (one for each agent) is denoted by $E$. The set of all possible collections of evidence for agent $i$ is denoted by $\mathbb{E}_{i}$ and we denote by $\mathbb{E}$ ($=\Pi_{i\in\mathcal{I}}\mathbb{E}_{i}$), the set of all possible profiles of evidence collections. As in the illustrating example, each collection of evidence can be associated with a set of subsets of the state space - namely those subsets in which the article occurs with positive probability.
We also assume that it is possible for an agent to submit the entire collection of evidence she is endowed with. This condition, which is commonly called normality, is present in Ben-Porath and Lipman (2012) as well, and is, generally speaking, a common feature of single stage mechanisms in such settings. [^9] Kartik and Tercieux (2012) is a notable exception. In case of uncertain evidence though, there emerges a subtle difference. In Ben-Porath and Lipman (2012), normality is formulated by requiring that a single “most informative” evidence message be available in any state to any agent. [^10] In context of the example for instance, the article $\{M,H\}$ is more informative than the article $\{L,M,H\}$. In general, the most informative article of evidence is given by the intersection of all the available articles. In this setting however, it is essential that agents be allowed to submit an entire collection of evidence, rather than just the most informative message. Intuitively, several different evidence collections can produce identical “most informative” messages, though these collections might be distinguishable when the designer observes each piece individually. See online appendix for formal details.
Evidence at state $s$ is distributed according to the commonly known prior $p:S\rightarrow\Pi_{i\in\mathcal{I}}\Delta(\mathbb{E}_{i})$. Denote the marginals of $p(s)$ on $\mathbb{E}_{i}$ and $\Pi_{j\neq i}\mathbb{E}_{j}$ by $p_{i}(s)$ and $p_{-i}(s)$, respectively. Deterministic evidence is a special case of this (the above distribution is degenerate). Note that $p$ is defined as a product distribution, eschewing the possibility of the evidence draws being correlated. This assumption is relaxed in Section 6.
Let $P_{i}$ denote the set of all possible evidence distributions for agent $i$, i.e. $P_{i}=\{p_{i}(s):s\in S\}$. We denote by $P$, the set $\Pi_{i\in\mathcal{I}}P_{i}$.
3.2 Mechanisms and Implementation
Any mechanism $\mathcal{M}$ is represented by the tuple $(M,g,\tau)$ where $M$ is the message space, $g:M\rightarrow A$ is the outcome function, and $\tau=\left(\tau_{i}\right)_{i\in\mathcal{I}}$ is the profile of transfer rules with $\tau_{i}:M\rightarrow\mathbb{R}$. At state $s$, and with a prior $p$, a mechanism $\mathcal{M}$ induces a Bayesian game $G(\mathcal{M},u,s,p)=\left\langle\mathcal{I},s,p,\left\langle M_{i},u_{i},S\times\mathbb{E}_{i}\right\rangle_{i\in\mathcal{I}}\right\rangle$ with the following properties:
  • The type space of a typical agent $i$ is: $S\times\mathbb{E}_{i}$.
  • Every type of every agent forms beliefs on $\mathbb{E}_{-i}$ according to $p_{-i}$.
  • The action space of a type $(s,E_{i})$ is given by $M_{i}(E_{i})$ (as some messages may be contingent on evidence availability). We only consider finite mechanisms, so that $M_{i}(E_{i})$ is finite for any $E_{i}$.
  • A typical (mixed) strategy for an agent $i$ at the state $s$ is a function: $\sigma_{i}:S\times\mathbb{E}_{i}\rightarrow\Delta(M_{i})$ such that $Supp(\sigma_{i}(s,E_{i}))\subseteq M_{i}(E_{i})$.
The common prior type space is specified by $p$. Given a game $G(\mathcal{M},u,s,p)$, say $\sigma_{i}(s,\cdot)$ is a best response to $\sigma_{-i}(s,.)$ if
$\sigma_{i}(s,E_{i})\left(m_{i}\right)>0\Rightarrow$
$m_{i}\in\arg\max_{m_{i}^{\prime}\in M_{i}}\sum_{E_{-i}\in\mathbb{E}_{-i}}p_{-i}(s)(E_{-i})\sum_{m_{-i}\in M_{-i}}\sigma_{-i}(s,E_{-i})(m_{-i})[u_{i}(g(m_{i}^{\prime},m_{-i}),\tau_{i}(m_{i}^{\prime},m_{-i}),s)].$
Next, we define the notion of a Bayesian Nash equilibrium in this particular setting.
Definition 1
A Bayesian Nash equilibrium (henceforth BNE) of the game $G(\mathcal{M},u,s,p)$ is defined as a profile of strategies $\sigma$ such that $\forall i,s$, $\sigma_{i}(s,\cdot)$ is a best response to $\sigma_{-i}(s,.)$.
A pure-strategy BNE is a BNE $\sigma$ such that for any $i$ and $s$, $\sigma_{i}\left(s,E_{i}\right)\left(m_{i}\right)=1$ for some $m_{i}$. We then define the main notion of implementation we work with in this paper.
Definition 2
An SCF $f$ is implementable in mixed-strategy BNE by a finite mechanism if there exists a finite mechanism $\mathcal{M}=(M,g,\tau)$ such that for any profile of bounded utility functions $\bar{u}$, at any state $s$, we have $g(m)=f(s)$ and $\tau(m)=0$ for every message $m\in M$ in the support of $\sigma(s)$ of any BNE $\sigma$ of the Bayesian game $G(\mathcal{M},u,s,p)$.
That is, for any BNE of the Bayesian game induced by the mechanism, the outcome is correct, and there are no transfers. We stress here that implementation should obtain regardless of the realized profile of evidence collections. Since we wish to implement by depending solely on evidence, rather than depending on preference variation, the notion of implementation above requires that implementation obtain regardless of the profile of bounded utility functions.
3.3 Lies and their classification
At the heart of the Bayesian implementation result is a classification of lies (states not identical to the truth), which we illustrate below.
Definition 3
A collection of evidence $E_{i}$ refutes a state $s$ (denoted $E_{i}\downarrow s$) if agent $i$ does not have $E_{i}$ or a superset of $E_{i}$ with positive probability at $s$. Otherwise, $E_{i}\not\downarrow s$.
That is, $E_{i}\downarrow s$ if $p_{i}(s)(E_{i}^{\prime})=0$ $\forall E_{i}^{\prime}\supseteq E_{i}$. In a mechanism, agents will always have the opportunity to withhold some articles from a collection. Therefore, even if a collection does not occur at $s$ with positive probability, being presented with this collection does not directly imply that $s$ is false. Indeed the evidence presented may be part of a larger collection which can occur. Therefore, only when a collection or any superset does not occur at $s$, can we claim that it refutes $s$.
In the context of the leading example with complete information, in state $hhh$ or $mmm$, the article $\{m,h\}$ refutes the state $lll$. A difference from the deterministic evidence setting is that it is not available with probability 1. For instance, if some firm claims that another firm’s signal is $l$ (when it is actually $m$), the latter firm cannot necessarily prove this claim to be false by presenting the article $\{m,h\}$ because it might be unavailable. This is different from a setting of deterministic evidence, where the necessary condition (measurability) entails that the realization of the evidence vary between states, so that the article $\{m,h\}$ would have to be available with probability 1.111111In a setting of complete information (e.g. Ben-Porath and Lipman (2012)) , an SCF $f$ satisfies measurability if whenever $f(s)\neq f(s^{\prime})$, there is an agent $i$ whose evidence endowment varies between $s$ and $s^{\prime}$.
Also, if a firm, when possessed with the evidence $\{h\}$ claims to have received the signal $m$, then it may be the only agent who can prove this claim to be false. [^12] Note that in context of this example, it might not be to the advantage of this firm to make such a claim, but since we do not rely on preference variation, in the general case, such issues may easily arise. This scenario, referred to as a self-refutable lie, also exists under deterministic evidence, but under deterministic evidence, the firm would be able to refute a lie about itself with probability 1, rather than being able to do so only with some probability (as in the above example).
The following conditions arise as a result of the definition of refutation used in our setup:
  1. 1.
    (se1) For any agent $i$,$\ p_{i}(s)(E_{i}\mathcal{)}=0$ if $E_{i}\downarrow s$. That is, at any state, the probability that collections of evidence which refute the truth are available is zero. In short, proof is true.
  2. 2.
    (se2) If $E_{i}\not\downarrow s^{\prime}$ then $p_{i}(s^{\prime})(E_{i}^{\prime})>0$ for some $E_{i}^{\prime}\supseteq E_{i}$. That is, if a collection of evidence $E_{i}$ does not refute $s^{\prime}$, then it (or a superset) is available in state $s^{\prime}$ as well.
In the online appendix, we illustrate the connections between this setup and the deterministic evidence setup from Ben-Porath and Lipman (2012).
Definition 4
A lie $s^{\prime}\in S$ is said to be refutable by $i$ at $s$ if $\exists E_{i}\in\mathbb{E}_{i}$ s.t. $p_{i}(s)(E_{i})>0$ and $E_{i}\downarrow s^{\prime}$.
That is, at $s$, $s^{\prime}$ is refutable by $i$ if $i$ might have a collection of evidence $E_{i}$ such that $E_{i}\downarrow s^{\prime}$. Clearly from (se1), $E_{i}$ is not available at $s^{\prime}$ (if it were available, then $E_{i}$ cannot refute $s^{\prime}$). Then, the support of the evidence distribution for agent $i$ contains a collection under the truth which is missing under a refutable lie.
Definition 5
A lie $s^{\prime}\in S$ is said to be refutable at $s$ if there is an agent $i$ so that $s^{\prime}$ is refutable by $i$ at $s$. Otherwise, it is nonrefutable.
Returning to the leading example, $hhh$ is nonrefutable at $mmm$. This brings up another difference from the deterministic evidence setting, which is that it is not possible for a firm to prove that its signal is $h$ rather than $m$. This is because there is no additional article of evidence available at state $hhh$ which is not available at state $mmm$. This scenario, called a nonrefutable lie (e.g. $hhh$ is nonrefutable at $mmm$), also occurs under deterministic evidence, but under deterministic evidence, it is necessary that some agent has additional evidence under a nonrefutable lie than under the truth with probability 1. Indeed the treatment of nonrefutable lies is one of the main differences between the deterministic and uncertain evidence setups.
We note that if $s^{\prime}$ is a nonrefutable lie at $s$, then no agent may possess (with positive probability) any articles of evidence in state $s$ which they may not possess at state $s^{\prime}$, since these articles would refute $s^{\prime}$. Therefore, if $s^{\prime}$ is nonrefutable at $s$, there are two possibilities with respect to the support of the evidence distribution at $s^{\prime}$. Either it strictly includes the support under $s$, so that $s$ is refutable (with positive probability) at $s^{\prime}$, or the supports are identical, so that only the probabilities with which various collections of evidence are available have changed. The latter scenario poses a major challenge in the implementation result we pursue.
4 Mixed-Strategy Bayesian Nash Implementation
4.1 A Necessary Condition: No Perfect Deceptions
Given that under deterministic evidence, the implementing condition (called measurability) entails only that at least some agent is endowed with a different set of evidence, it would be natural to expect that a stochastic version of this condition, which entails requiring $p(s)\neq p(s^{\prime})$ whenever $f(s)\neq f(s^{\prime})$ would suffice as the implementing condition for Bayesian implementation. We term this condition stochastic measurability. It turns out, however, that not every nonrefutable lie can be eliminated by a mechanism, necessitating a stronger implementing condition. We establish the condition and prove its necessity after some preliminaries below.
First, we define a deception for an agent of type $(s,E_{i})$.
Definition 6
A deception for an agent $i$ is a function $\alpha_{i}:S\times\mathbb{E}_{i}\rightarrow S\times\Delta\mathbb{E}_{i}$ such that if $(s^{\prime},E_{i}^{\prime})\in\operatorname{supp}\alpha_{i}(s,E_{i})$, then $E_{i}^{\prime}\subseteq E_{i}$.
We interpret $\alpha_{i}$ as being a function from the agent’s true type to a distribution over types she can mimic (which requires that true type has a weakly larger collection of evidence than the mimicked type). [^13] A critical difference between the notion of deception in Jackson (1991) and the one used here is that deceptions do not permit mixing in the state type. The mechanism we propose later will be able to detect and prevent such mixing, so that we need not allow for such deceptions in the formulation of the necessity condition. A profile of deceptions is denoted by $\alpha=(\alpha_{i})_{i\in\mathcal{I}}$. This notion is similar to the one with the same name in Jackson (1991).
Note that a deception can also be interpreted as a strategy for an agent in a direct mechanism (a mechanism where agents only report their types, i.e. $M_{i}=S\times\mathbb{E}_{i}$ for each agent).
Definition 7
A deception $\alpha_{i}$ is perfect for state $s^{\prime}$ at state $s$ for agent $i$ if,
$\forall E_{i}^{\prime}\text{,}\sum_{E_{i}\in\mathbb{E}_{i}}p_{i}(s)(E_{i})\alpha_{i}(s,E_{i})(s^{\prime},E_{i}^{\prime})=p_{i}(s^{\prime})(E_{i}^{\prime})\text{.}$
Likewise, a profile of deceptions $\alpha=(\alpha_{i})_{i\in\mathcal{I}}$ is perfect for state $s^{\prime}$ at state $s$ if $\alpha_{i}$ is perfect for state $s^{\prime}$ at state $s$ for each agent $i$.
That is, a deception is perfect for state $s^{\prime}$ at state $s$ if every agent induces the same distribution over her types as she would if the true state were $s^{\prime}$ and if she were reporting truthfully.
We denote by $s\rightarrow_{i}s^{\prime}$ the notion that agent $i$ has a perfect deception from $s$ to $s^{\prime}$. Therefore, if there is a perfect deception from $s$ to $s^{\prime}$, then, $s\rightarrow_{i}s^{\prime}$ $\forall i\in\mathcal{I}$. Further, $s\not\rightarrow_{i}s^{\prime}$ denotes the notion that $i$ does not have a perfect deception from $s$ to $s^{\prime}$ and $s\not\rightarrow s^{\prime}$ denotes that there is no perfect deception from $s$ to $s^{\prime}$.
As mentioned earlier, it is not possible to eliminate nonrefutable lies such that there is a perfect deception for the nonrefutable lie at the true state. We now define the necessary condition for implementation.
Definition 8
(Condition NPD - No Perfect Deceptions) We say that a social choice function $f$ satisfies NPD whenever for two states $s$ and $s^{\prime}$, if at $s$, state $s^{\prime}$ is nonrefutable, and there exists a perfect deception for $s^{\prime}$, then $f(s)=f(s^{\prime})$.
That is, to have different desirable outcomes between $s$ and $s^{\prime}$, either more evidence must be available at $s$ than is available at $s^{\prime}$, or it must be impossible to have a perfect deception for state $s^{\prime}$ at state $s$. We note that one way for a perfect deception to be impossible is if there are articles of evidence at $s^{\prime}$ which are not available at $s$. We can therefore also interpret the above condition as follows: we are able to eliminate all kinds of lies other than nonrefutable lies so that agents can mimic the lie both in the state and the evidence dimensions perfectly. [^14] For instance, if any pair of states $s$ and $s^{\prime}$ are mutually refutable, then any SCF satisfies NPD. Further see online appendix Section 1 for a discussion of possible cases of refutation and perfect deceptions for a pair of states $s$ and $s^{\prime}$.
We also wish to stress the fact that the existence of a perfect deception from $s$ to $s^{\prime}$ does not by itself preclude our ability to separate the states by a mechanism. Indeed it must be the case that $s^{\prime}$ is nonrefutable at $s$. To see this, consider a scenario where the evidence structure is identical between $s$ and $s^{\prime}$ except that each collection at $s$ is appended with an additional article of evidence $e_{i}$ for each agent $i$. Clearly, a perfect deception involves each agent claiming the state is $s^{\prime}$ and withholding the article $e_{i}$. However, it is clear that these states can be separated by incentivizing refutation of a state claim by presenting $e_{i}$. [^15] For instance, this is true of the motivating example with the perturbation including the article of evidence $\{h\}$ at the state $hhh$.
For an example of a social choice function and an evidence structure under which the evidence distribution varies between two states, we refer the reader to the leading example in Section 2. We reproduce the evidence structure below:
Signal Received Evidence Distribution
$h$ $0.6\times\{\{m,h\},\{l,m,h\}\}+0.4\times\{\{l,m,h\}\}$
$m$ $0.4\times\{\{m,h\},\{l,m,h\}\}+0.6\times\{\{l,m,h\}\}$
$l$$\{\{l,m,h\}\}$
It is clear that the distribution of evidence does change between every pair of states, so that every SCF is stochastically measurable. However, it can be seen that it is possible to mimic the distribution at $mmm$ when the true state is $hhh$, as the article $\{m,h\}$ is available with extra probability at $hhh$. For instance, the deception $\alpha_{A}(h,\{\{m,h\},\{l,m,h\}\})\rightarrow\frac{1}{3}(m,\{\{l,m,h\}\})+\frac{2}{3}(m,\{\{m,h\},\{l,m,h\}\})$ suffices to do so. However, it is not possible to mimic the distribution at $hhh$ when the state is $mmm$, as the article $\{m,h\}$ is not available often enough. This establishes an important property of the NPD condition, that is, it is not bidirectional. Therefore, the structure induced by the social choice function on the state space is not partitional. We stress here that this is a stark point of differentiation between deterministic and uncertain evidence, since the necessary condition under deterministic evidence - measurability, induces a partitional structure on the state space.
We present the main result of this section below.
Theorem 1
A social choice function is implementable in BNE if and only if it satisfies NPD.
We present the proof of necessity below, while the implementing mechanism which constitutes the sufficiency proof is presented in Section 4.4.
Proof. (Necessity of NPD) Suppose a mechanism $\mathcal{M}=(M,g,\tau)$ implements $f$ and consider a pair of states $s$ and $s^{\prime}$ so that $s^{\prime}$ is nonrefutable at $s$ and there is a perfect deception from $s$ to $s^{\prime}$.
First, we claim that
This follows from the nonrefutability of $s^{\prime}$ at $s$.
Now, consider any equilibrium $\sigma$ of the mechanism and the following strategy for agent $i$ at state $s$ when endowed with the collection $E_{i}$. First, the agent plays according to the deception $\alpha_{i}(s,E_{i})$. This yields another type realization, which we denote by $(s^{\prime},E_{i}^{\prime})$. Now, the agent plays $\sigma_{i}(s^{\prime},E_{i}^{\prime})$. Since there is no additional collection of evidence at $s$, this object is well defined. With some abuse of notation, we denote the profile of such strategies by $\sigma\circ\alpha$.
We claim that $\sigma\circ\alpha$ forms an equilibrium at $s$ with outcome $f(s^{\prime})$. To see this, consider an arbitrary agent $i$. First, given the strategies of other agents $j\neq i$, the belief induced over the actions of other agents $j$ is the same under $\sigma\circ\alpha$ at $s$ as that under $\sigma$ at the state $s^{\prime}$. Given these beliefs, at state $s^{\prime}$, $\sigma$ induces the outcome $f(s^{\prime})$ with no transfers. Each type of agent $i$ can obtain this outcome in state $s$ by playing according to $\sigma_{i}\circ\alpha_{i}$. If any type $E_{i}$ of agent $i$ at $s$ has a profitable deviation, then there is another type of agent $i$ at $s^{\prime}$ which has weakly more evidence than $E_{i}$ (from the claim (LABEL:*) above) and can also deviate profitably. This contradicts the optimality of $\sigma_{i}$ against $\sigma_{-i}$ at $s^{\prime}$. Therefore, playing according to $\sigma_{i}\circ\alpha_{i}$ continues to be a best response for the type. Thus, if there is a perfect deception $\alpha$ from $s$ to $s^{\prime}$, then any implementable social choice function must have $f(s)=f(s^{\prime})$.
Informally, under $\alpha$, agents pretend to be other types and play the equilibrium strategy under $\sigma$ of the mimicked type. The reason a type is best responding even though she may be playing the equilibrium strategy of another type is that the outcome is the same and the preferences of agents do not change from state to state. That is, even though the concerned type may have extra evidence, and therefore be able to play other messages, there was another type at $s^{\prime}$ who also had that extra evidence and that type’s best response still led to the outcome $f(s^{\prime})$ which this agent is also getting from her strategy under $\sigma\circ\alpha(s)$, so that she continues to best respond. Therefore, since an equilibrium outcome at $s$ is $f(s^{\prime})$, then $f$ is implementable only if $f(s)=f(s^{\prime})$.
We note here that whereas the discussion above has centered around deceptions where an agent pretends to mimic herself at another state, this applies to mimicking a profile of distributions as well. More precisely, suppose that in the true state $s^{\ast}$, which is characterized by the profile of distributions $(p_{i}^{\ast})_{i\in\mathcal{I}}$, agents pretend to mimic themselves at another state $s^{\prime}$, which is characterized by a profile of distributions $(p_{i}^{\prime})_{i\in\mathcal{I}}$. If they can perfectly mimic this other state (in terms of playing their equilibrium strategies for state $s^{\prime}$), then by Theorem 1, $f$ is implementable only if $f(s^{\ast})=f(s^{\prime})$.
This yields an interesting insight: since it is always possible to mimic any cheap talk messages in the message space, nonrefutable lies $s^{\prime}$ for which agents do not have perfect deceptions from the truth must require presentation of evidence which is impossible in the true state. Then, if an agent $i$ does not have a perfect deception at state $s$ for $s^{\prime}$, then she cannot mimic the distribution $p_{i}^{{}^{\prime}}$, as otherwise it would be possible for her to play the equilibrium strategies for the state $s^{\prime}$. This highlights the role of evidence in this setup, which is twofold. Either agents can refute distributional claims using evidence, or agents are unable to mimic incorrect states perfectly owing to their inability to mimic the distribution of evidence.
4.2 A Challenge Scheme
The above discussion established that when agents report certain nonrefutable lies, they are unable to mimic the distribution of evidence for these lies. This brings up the question of how we can exploit this fact in the construction of mechanisms to facilitate the elimination of nonrefutable lies. To elucidate the underlying idea, we first consider an analogous situation.
Consider a setting with two agents. The first agent has a deck of cards, from which she draws a card and presents it. The second agent knows that the first’s deck of cards is missing one black card. How does she convey this information credibly to the designer if it is not possible to check the entire deck? If she simply says there is a card missing, this claim may not be credible. Rather, the second agent can place a bet of the following form - if you draw a black card, take a dollar from me. If you draw a red card, give me 99 cents. This bet clearly loses the second agent money if there is no card missing. If she believes there is a black card missing though, then this wins her money, as $\frac{26}{51}\times 0.99+\frac{25}{51}\times(-1)>0$. The fact that the second agent is willing to take this bet credibly signals to the designer that there is something wrong with the deck.
Note that the deck of cards is analogous to agent 1’s mixed-strategy. Therefore, it is impossible to check the deck of cards and the second agent must find this alternate solution to inform the designer that the first agent is not playing the truthtelling strategy. Individual cards are analogous to collections of evidence, and the inability to induce the correct distribution of evidence is analogous to not having a black card. For the implementation problem then, agents can blow the whistle on an incorrect (nonrefutable) state report by identifying which agent does not have a perfect deception from the truth to the nonrefutable lie, and then betting sums of money on that agent’s plausible collections of evidence so that - (i) If the nonrefutable lie were to be the truth, then the bet loses money, and (ii) Since this is not the case, the bet actually wins money. As we will discuss later, this can be thought of as analogous to a monotonicity like condition, yielding a “reversal” between true and false states. Returning to the illustrating example, if a firm tried to claim $hhh$ when the true state was $mmm$, it would, at best, be able to induce probability $0.4$ on the article $\{m,h\}$ as that is the maximum probability with which it is available in state $mmm$. Then, consider the following bet: $1$ dollar on $\{\{l,m,h\}\}$ and $1$ dollar against $\{\{l,m,h\},\{m,h\}\}$. In state $lll$ under truthful reporting, it yields an expected loss of 20 cents, while in state $mmm$, it yields a minimum expected profit of 20 cents. We will show below that this is possible whenever there is no perfect deception. Before that, we will provide a simple way for an agent to place such a bet.
Asking agents to place bets on (say) agent $i$’s evidence would mean that agents would have to provide the designer a vector of numbers (one for each collection in $\mathbb{E}_{i}$). This yields a complex message space. In fact, the designer can make the appropriate bets for agents. All that she needs to know is what state (say $s^{\prime}$) is being claimed, and what is the true state. With this information, the designer can deduce which agent has no perfect deception for the claimed state and place an appropriate bet against the evidence of that agent on behalf of the agent who wishes to raise the challenge. We will denote these bets by $b_{i}:S\times S\rightarrow\mathbb{R}^{|\mathbb{E}_{i}|}$.
Remark 1
The above can also be viewed as the building block of a monotonicity like condition. That is, for a pair of states $s$ and $s^{\prime}$ such that $s\not\rightarrow_{i}s^{\prime}$, if all agents claim that the state is $s^{\prime}$, then there is an agent $j$ for whom $(f(s^{\prime})+b_{\hat{\imath}}(s^{\prime},s)(E_{\hat{\imath}}))$ is strictly preferred to $f(s^{\prime})$ in state $s$ while $f(s^{\prime})$ is weakly preferred to $(f(s^{\prime})+b_{\hat{\imath}}(s^{\prime},s)(E_{\hat{\imath}}))$ in state $s^{\prime}$. The allocation $f(s^{\prime})+b_{\hat{\imath}}(s^{\prime},s)(E_{\hat{\imath}})$ therefore, forms a test allocation which yields a reversal for an agent $j$ with $f(s^{\prime})$ between the states $s$ and $s^{\prime}$.
We formalize this notion below. First, denote by $\mathcal{A}_{i}(s,E_{i})$ the set of all deceptions $\alpha_{i}(s,E_{i})$ for agent $i$ when endowed with evidence $E_{i}$ at state $s$. Let $\alpha_{i}(s)$ be defined as $\alpha_{i}(s)=(\alpha_{i}(s,E_{i}))_{E_{i}^{\prime}\in\mathbb{E}_{i}}$, and analogously, let $\mathcal{A}_{i}(s)$ be the set of all possible profiles $\alpha_{i}(s)$. Further, denote by $p_{\alpha_{i},s}(E_{i}^{\prime})$ the probability induced over evidence collection $E_{i}^{\prime}$ by agent $i$ under $\alpha_{i}(s)$. That is,
$p_{\alpha_{i},s}(E_{i}^{\prime})=\sum_{E_{i}\in\mathbb{E}_{i}}p_{i}(s)(E_{i})\alpha_{i}(s,E_{i})(s^{\prime},E_{i}^{\prime})\text{.}$
Defining probability distributions over evidence, i.e. $(p_{\alpha_{i},s}(E_{i}^{\prime}))_{E_{i}^{\prime}\in\mathbb{E}_{i}}$by $p_{\alpha_{i},s}$, we then denote by $P_{i}(s)$ the set $\{p_{\alpha_{i},s}:\alpha_{i}\in\mathcal{A}_{i}\}$, which is the set of all possible such probability distributions which agent $i$ can induce by playing any deception in state $s$. Among these distributions is also $p_{\alpha_{i}^{\ast},s}$, the distribution over agent $i$’s evidence at $s$, which corresponds to the “no deception” scenario. At state $s^{\prime}$ let $p_{\alpha_{i}^{\ast},s^{\prime}}$ denote the induced distribution over evidence by the “no deception” scenario.
With these preliminaries established, we now present the challenge scheme in the following lemma.
Lemma 1
For every agent $i$, there is a finite set $B_{i}=\{b_{i}(s,s^{\prime}):s\not\rightarrow_{i}s^{\prime}\}$ such that for any pair of states $s$ and $s^{\prime}$ for which $s\not\rightarrow_{i}s^{\prime}$, we have $b_{i}(s,s^{\prime})\cdot p_{\alpha_{i}^{\ast},s^{\prime}}<0$ and $b_{i}(s,s^{\prime})\cdot p>0$ $\forall p\in P_{i}(s)$.
Proof. Since $s\not\rightarrow_{i}s^{\prime}$, $p_{\alpha_{i}^{\ast},s^{\prime}}\notin P_{i}(s)$. Further, $P_{i}(s)$ is closed and convex. Then, from the separating hyperplane theorem, there is a hyperplane $b_{i}$ such that $b_{i}\cdot p_{\alpha_{i}^{\ast},s^{\prime}}<0$ and $b_{i}\cdot p>0$ $\forall p\in P_{i}(s)$.
For a pair of states $s$ and $s^{\prime}$ one such $b_{i}$ is sufficient. Iterating over all possible pairs of states which satisfy the premise of no perfect deception yields the set $B_{i}$. Therefore, each $B_{i}$ is finite.
Remark 2
The above lemma is interpreted as follows: there is a vector of real numbers (one for each possible collection of evidence for $i$) so that under truthful revelation in state $s^{\prime}$, the expectation of this vector is negative whereas for any possible strategy of agent $i$ in state $s$, the expectation is positive. Accordingly, if an agent $j$ were to receive/give sums of money based on $b_{i}$ according to agent $i$’s evidence presentation in the direct mechanism, then in state $s^{\prime}$ under truthful revelation by agent $i$, agent $j$ loses money while under any strategy for agent $i$ in state $s$, agent $j$ gains money. This yields a monotonicity like interpretation - if $s^{\prime}$ is a lie, the agent $j$ can ”blow the whistle” and indeed chooses to do so since it is profitable.
We remark here that the set $B_{i}$ is a set of possible bets, one for each pair of states such that there is no perfect deception for one to another. For the set to be usable in a mechanism, we need a way for an agent to utilize members of this set to challenge unanimous but incorrect state claims. This operationalization is achieved as follows. First, define a function $\hat{\imath}:S\times S\rightarrow\mathcal{I}$ such that if $s\not\rightarrow s^{\prime}$, $s\not\rightarrow_{\hat{\imath}(s,s^{\prime})}s^{\prime}$. That is, $\hat{\imath}$ identifies the agent who does not have a perfect deception from state $s$ to $s^{\prime}$. Then, $b_{\hat{\imath}}:S\times S\rightarrow B_{\hat{\imath}}$ is defined such that $b_{\hat{\imath}}(s,s^{\prime})\cdot p_{\sigma_{\hat{\imath}}^{\ast},s^{\prime}}<0$ and $b_{\hat{\imath}}(s,s^{\prime})\cdot p>0$ $\forall p\in P_{\hat{\imath}}(s)$. The existence of such functions follows immediately from Lemma 1.
That is, when an agent $j$ wants to bet against a report of $s^{\prime}$ in state $s$, then merely telling the designer that the true state is $s$ suffices. This is because the designer can use the function $\hat{\imath}$ as defined above to infer which agent has no perfect deception from $s$ to $s^{\prime}$ and can appropriately place a bet against $\hat{\imath}$’s evidence on behalf of $j$ using the function $b_{\hat{\imath}}$ so that the bet loses money for $j$ if $s^{\prime}$ were indeed true and $\hat{\imath}$ were reporting truthfully, and wins money for $j$ if $s$ were true irrespective of what strategy $\hat{\imath}$ employs.
4.3 A Sketch of the Proof
In Section 4.1, we have established that NPD is necessary for implementation. In what follows, we will construct an implementing mechanism to establish that it is also sufficient. To fix ideas, we begin with a sketch of the proof.
Figure 1: An illustration of the mechanism
The mechanism asks each agent for the following reports: A distribution report about herself, a distribution report for the agent to her right, an evidence report, and a state report (used for placing bets). The proof utilizes the following steps. The first step is an intermediate result that is used frequently, and thus forms one of the main building blocks of the proof. In essence, the result states that if each type of a particular agent (say $i$) submits all the evidence it is endowed with, then every agent ($i$ herself, and all others) present truthful reports about $i$’s evidence distribution. This result is obtained as follows. First, we use a proper scoring rule ($\tau^{2}$ in the proof) to incentivize other agents to predict the evidence report by agent $i$. Since agent $i$ presents maximal evidence in all her types, the optimal predictor for other agents is the true distribution for agent $i$. This truthful report is used (by means of a crosschecking penalty) to discipline agent $i$’s report about herself. This is denoted by $\tau^{3}$ in the proof. To reiterate, agent $i$’s maximal evidence presentation incentivizes other agents to make truthful predictions about her, whereupon a crosscheck incentivizes agent $i$ to make truthful reports about her own distribution. A direct corollary of this result is that if all agents present maximal evidence, then the entire profile of reports is truthful.
A key novelty of this argument is in the solution to the central tension when using scoring rules - on the one hand, the scoring rule must be active on all realizations of evidence to elicit the true distribution (otherwise it elicits the distribution conditional on being active, which may not yield truthful elicitation), but on the other hand, we require zero transfers on equilibrium. This is solved by breaking the scoring rule transfer into two parts - a direct part, which incentivizes predictions, and an offset part, which is used to cancel the direct part in equilibrium. When an agent makes a prediction, she only controls the direct part, so that from her point of view, the scoring rule is always active. This incentivizes her to make truthful predictions about others. However, her truthful report is used to discipline other agent’s reports about themselves, and this disciplined report is used to offset the scoring rule, so that it does not result in a transfer on equilibrium.
Armed with this intermediate result, the second step argues that the report profile must be pure (recall that agents are allowed to mix) and must be consistent with some state $s$. This step depends on an incentive for evidence presentation which is triggered if agents disagree with each other. This is denoted by $\tau^{1}$ in the proof. Broadly, if agents mix, then agents expect disagreement in the distribution reports, and if the message profile is inconsistent (due to such disagreements), then each agent is incentivized to present their maximal evidence. Then, from the above intermediate result, we can argue that the message profile cannot be inconsistent, since agents must make truthful reports if all evidence presentation is maximal and truthful reports must, by definition, be consistent with the true state. Figure 1 presents an overview of this step to help fix ideas.
The third step establishes that there cannot be a consensus on a refutable lie. We institute a penalty ($\tau^{4}$ in the proof) which penalizes agents if this commonly agreed upon state $s$ is refuted by another agent. There is also a small reward for refuting this consensus (this is implicit within $\tau^{1}$, which is also activated upon refutation). This penalty ($\tau^{4}$) is chosen to be large enough to deter agents from agreeing on a refutable lie.
The fourth and final step establishes that there cannot be a consensus on a nonrefutable lie which has a different outcome than the truth. If this were to be the case, then from Condition NPD, there is an agent (say $i$) who does not have a perfect deception from the true state to $s$. Then, the mechanism provides a way for agents to make a bet, as discussed in Section 4.2. Further $\tau^{1}$ is activated when there is an active bet as well. This leads to the submission of maximal evidence by all agents, which by the first step means that the profile of reports must have been true, so that there could not have been a consensus on a nonrefutable lie in the first place.
Thus, we are left only the possibility of a joint pure report of the true state (or a state which is outcome equivalent to the truth), so that implementation obtains. In Section 4.5.5, we also show that there are no transfers in equilibrium.
4.4 Mechanism
4.4.1 Message Space
Suppose the agents are seated around a circular table, so that for each agent $i$, there is an agent to her right (this is agent $i+1$). For agent $I$, the agent considered as “being to the right of” her is agent $1$. With this, the message space is defined as follows,
$M_{i}=P_{i}\times P_{i+1}\times\mathbb{E}_{i}\times S\text{,}$
and a typical message ($m_{i}$) is given by $m_{i}=(p_{i,i},p_{i,i+1},E_{i},s_{i})$.
That is, each agent makes a claim about her own evidence distribution ($p_{i,i}$) , a claim about the evidence distribution of the agent to her right ($p_{i,i+1}$), submits a collection of evidence and makes a state claim for (possibly) betting against other’s evidence reports.
4.4.2 Outcome
A message profile $m$ is said to be consistent with state $s$ if each distributional claim matches the true distribution in state $s$. That is, $m=(p_{i,i},p_{i,i+1},E_{i},s_{i})_{i\in\mathcal{I}}$ is consistent with $s$ if $\forall i,j$, $p_{i,j}=p_{j}(s)=p_{j,j}$. Otherwise, it is inconsistent.
The outcome is given by $f(s)$ if $m$ is consistent with a state $s$, and an arbitrary outcome $a\in A$ otherwise.
4.4.3 Transfers
The mechanism uses five transfers. There are scaling parameters with most of the transfers, and we will define the appropriate scaling in Section 4.4.4.
The first transfer provides an agent the incentive to submit evidence when there is inconsistency, or an active bet, or if a consistent profile has been refuted by an agent. That is,
[…]
where $\varepsilon>0$ is a small positive number.
For defining the second transfer (which is a modified proper scoring rule), we first define a general scoring rule which will find use in the rest of this paper. A scoring rule requires two inputs to score a prediction - (i) A claimed distribution (say $p$) over the possible realizations of a random variable (say $X$); and (ii) The actual realization (say $x$). Accordingly, $Q(p,x)$ is defined as $[2p(x)-p\cdot p]$. It is well known that this scoring rule elicits truthful belief elicitation. With this, define
$\tau_{i,i+1}^{2}(m)=\left\{\underline{\tau}\times[Q(p_{i,i+1},E_{i+1})-Q(p_{i+1,i+1},E_{i+1})]\text{,}\right.$
Where $\underline{\tau}>0$. This transfer provides agent $i$ the incentive to truthfully report agent $i+1$’s distribution (due to the first term) when agent $i+1$ presents all her evidence. In equilibrium, the reports are consistent and therefore the transfer is off.
[…]
Every agent’s report about herself is cross-checked with the report about her by the agent to her left. If the reports are not in agreement, then the agent is fined $\underline{\tau}$.
The fourth transfer is related to lies which other agents can refute.
[…]
That is, an agent $i$ is fined $\overline{\tau}$ if another agent has refuted state $s$ obtained from the vector of reports by agents about the agents to their right. Note that this transfer does not apply to an agent’s report about herself. Further, this transfer is inactive when the profile $(p_{i,i+1})_{i\in\mathcal{I}}$ is not consistent.
$\tau^{5}$ is as defined below.
[…]
where $\varepsilon>0$ is a small positive number as defined earlier. Note that $\tau^{1}$ and $\tau^{5}$ are scaled using the same number $\varepsilon$.
That is, if agent $i$ bets against a unanimous report consistent with state $s^{\prime}$ with a bet claiming state $s_{i}$, the designer evaluates the challenge in accordance with the set $B_{j}$, where $j=\hat{\imath}(s^{\prime},s_{i})$ and the function $b_{j}$ defined in Section 4.2 and pays out the revenue from the bet to agent $i$.
4.4.4 Scaling
In this section, we will define some parameters used for appropriately scaling the transfers, and then establish that there are values for $\overline{\tau}$, $\varepsilon$, and $\underline{\tau}$ so that the appropriate scaling is possible.
First, we provide an intuitive overview of the scaling based on the proof sketch in Section 4.3. Recall that we are working in a normalized setup in which the utility of the outcome has been normalized to a number less than 1. Since an arbitrarily small $\varepsilon$ provides sufficient incentive for agents to submit evidence and make bets, we can choose $\varepsilon$ small enough so that a penalty of 1 dollar is enough to dominate not only any change in outcome, but also any gains or losses from $\tau^{1}$ or $\tau^{5}$. The scoring rule $\tau^{2}$ is at the medium scale ($\underline{\tau}$) and dominates any effect from a change of outcome. [^16] This is, strictly speaking, larger than it needs to be, since $\tau^{2}$ needs only to dominate the transfers at $\varepsilon$ scale (as in setting her report about the agent to her right, an agent may cause (or break) consistency). However, we scale it at the level of $\underline{\tau}$ for notational parsimony. The crosscheck $\tau^{3}$ is also at this scale since similar considerations apply to the agent’s report about herself. Since we wish to prioritize the elimination of lies which others can refute above other incentive considerations, $\tau^{4}$ dominates all other transfers, so that it belongs to the highest transfer scale ($\overline{\tau}$). We define the transfer scales formally below.
Since $\tau^{2}$ is a proper scoring rule, if an agent $i+1$ presents all her evidence, agent $i$ maximizes her revenues from $\tau^{2}$ by truthfully reporting agent $i+1$’s evidence distribution. We require that the minimum expected loss for an agent $i$ from lying about agent $i+1$’s distribution dominates any effect from a change in the outcome. That is,
$\min_{s\in S}\min_{p_{i,j}\neq p_{j}(s)}\sum_{E_{j}\in\mathbb{E}_{j}}p_{j}(s)(E_{j})\underline{\tau}[Q_{j}(p_{j}(s),E_{j})-Q_{j}(p_{i,j},E_{j})]>1$ (2)
Turning to the crosscheck, it must dominate any incentive from the outcome. Since the minimum loss from lying about agent $i+1$’s distribution scales linearly in $\underline{\tau}$, we can choose $\underline{\tau}>1$ such that inequality 2 is satisfied.
Further, define as $\bar{\tau}^{2}$ the maximum possible change in utility from $\tau^{2}$. That is, i.e.
$\bar{\tau}^{2}=\max_{i\in\mathcal{I}}\max_{E_{i+1}\in\mathbb{E}_{i+1}}\max_{p_{i,i+1},p_{i,i+1}^{\prime}\in P_{i+1}}\underline{\tau}[Q_{i+1}(p_{i,i+1},E_{i+1})-Q_{i+1}(p_{i,i+1}^{\prime},E_{i+1})]$
Now we define the fourth scaling inequality as follows:
$\overline{\tau}\geq\frac{1+\bar{\tau}^{2}}{\min_{s^{\prime}\in RL_{i}(s)}\{p_{i}(s)(E_{i}):E_{i}\downarrow s^{\prime}\}},\forall s\in S,\forall i\in\mathcal{I}\text{.}$ (3)
That is, $\tau^{4}$ provides sufficient incentive to avoid being caught lying even for the smallest probability with which refutable lies may be refuted by the evidence of other agents.
It is then clear that the parameters $\varepsilon$, $\underline{\tau}$ and $\overline{\tau}$ can be chosen to satisfy the above inequalities simultaneously.
4.5 Proof of Implementation
In what follows, we assume that the true state is $s^{\ast}$ and that it induces the profile of evidence distributions $(p_{i}^{\ast})_{i\in\mathcal{I}}$.
4.5.1 Preliminary Results
Lemma 2
If agent $i$ presents all her evidence in $E_{i}$, then each agent $j\in\mathcal{I}$ presents the truth in each message $p_{j,i}$.
Proof. If agent $i$ presents all her evidence in $E_{i}$, then for agent $i-1$, setting $p_{i-1,i}=p_{i}^{\ast}$ maximizes her payoff from $\tau^{2}$ (since truthful reporting is optimal under a scoring rule). However, this could also cause losses from $\tau^{1}$ (by creating consistency) and $\tau^{5}$ (by activating a bet via creating consistency). However, since $\underline{\tau}>1$ (by choice), $p_{i-1,i}=p_{i}^{\ast}$ is the best response for agent $i-1$ in equilibrium. Then, since $\underline{\tau}>1$, $p_{i,i}=p_{i}^{\ast}$ is the best response for agent $i$.
4.5.2 Consistency
Claim 1
In any equilibrium, there is a state $s$ such that the message profile is consistent with $s$.
Proof. If all agents present all their evidence, then Lemma 2 yields that all agents present true distributional claims in their reports and thus the profile of reports must be consistent with the truth, so that $s=s^{\ast}$.
Now, suppose that there is inconsistency but not all agents present all their evidence. We claim that this is only possible if precisely one agent $i$ randomizes in her reports $p_{i,i}$ or $p_{i,i+1}$. To see this, notice that if two or more agents randomize in their distribution reports, then each type of each agent expects inconsistency with positive probability and presents all their evidence. So suppose that only agent $i$ mixes. Then in each message, each agent $j\neq i$ expects inconsistency with positive probability and presents all her evidence ($\varepsilon>0$). Then, from Lemma 2, $p_{k,j}$ is the truth for all $k\in\mathcal{I}$ and $j\neq i$. Therefore, $i$ can only mix in $p_{i,i}$. However, by hypothesis, each agent $j\neq i$ presents a pure report in $p_{j,i}$. Therefore, in any message $m_{i}$ where $p_{i,i}\neq p_{j,i}$, she incurs a loss from $\tau^{3}$. Note that there are no losses from $\tau^{4}$ since agent $i$ is only changing $p_{i,i}$ and $\tau^{4}$ does not apply to an agent’s report about herself. Consider a deviation to match $p_{j,i}$. This avoids the loss from $\tau^{3}$, but may cause losses from $\tau^{1}$ (by creating consistency) and $\tau^{5}$ (by activating a bet via creating consistency). Since $\underline{\tau}>1$, this is a profitable deviation. Therefore, we have a contradiction.
4.5.3 Eliminating Refutable Lies
Claim 2
There is no equilibrium where the distribution reports are consistent with a refutable lie.
Proof. Suppose instead that $m$ is consistent with a refutable lie $s$. If it is refutable by two or more agents, then $\tau^{1}$ provides the incentive to refute the lie, and so all agents expect the lie to be refuted with positive probability so that they present maximal evidence. In turn, Lemma 2 implies the profile must be consistent with the truth, a contradiction.
So suppose $s$ can only be refuted by some agent $i$. Then, all agents other than $i$ expect agent $i$ to refute $s$ with positive probability and therefore present maximal evidence. Thus, from Lemma 2, $p_{.,j}=p_{j}^{\ast}$ $\forall j\neq i$. Then, $p_{i-1,i}\neq p_{i}^{\ast}$ (otherwise the profile would be consistent with the truth). Therefore, in presenting $p_{i-1,i}$, agent $i-1$ knows that this claim will be refuted with probability at least $\min_{E_{i}\in\mathbb{E}_{i}}\{p_{i}(s^{\ast})(E_{i}):E_{i}\downarrow s\}$. However, if she deviates to $p_{i-1,i}=p_{i}^{\ast}$, the profile has $p_{i,i+1}=p_{i+1}(s^{\ast})$ $\forall i$ and therefore cannot be refuted. So, the agent gains $\overline{\tau}$ from $\tau^{4}$. Also, the agent is deviating away from consistency, so that $\tau^{1}$ yields no losses. Since $\overline{\tau}$ dominates all the other incentives (Inequality (3)), deviating to $p_{i}^{\ast}$ instead of presenting $p_{i-1,i}$ is a profitable deviation for agent $i$.
4.5.4 No Bets in Equilibrium
Claim 3
In any equilibrium, the distribution reports are not consistent with a nonrefutable lie.
Proof. From Claim 1, the agents must all present pure distribution reports which are consistent with a state $s$. From Claim 2, $s$ is not refutable by any agent, since this would mean that an agent is presenting a lie which another agent can refute. If $f$ prescribes a different outcome at $s$ than at the truth, then there must be an agent $i$ such that $s^{\ast}\not\rightarrow_{i}s$, that is, agent $i$ does not have a perfect deception from the truth to $s$. Then, any agent $j\neq i$ has the incentive to bet with probability 1, since $\varepsilon>0$, whereupon all evidence presentation is maximal (due to $\tau^{1}$) and Lemma 2 presents a contradiction.
4.5.5 Implementation
Claim 4
In any equilibrium, the outcome is $f(s^{\ast})$ and there are no transfers.
Proof. From Claims 1-3, the report profile is consistent with the true state. Therefore, the outcome is $f(s^{\ast})$. Clearly, such a consensus is not refutable, since any collection available under the truth cannot (by definition) refute it.
Now, we claim that there are no bets in equilibrium. To see this, notice that if an agent $i$ bets in equilibrium, then it must be that agents have not presented all their evidence in $E$, since from Lemma 1, a bet yields a loss if placed against a truthful reporting strategy. If some agent $i$ finds it profitable to place a bet in any of her messages, then she must find it profitable to bet in all of her messages. Then, each agent finds it optimal to present all their evidence in $E$, a contradiction. Since the message profile is consistent, and features no bets in equilibrium, $\tau^{1}$, $\tau^{3}$ and $\tau^{5}$ are inactive as well.
Finally, notice that the absence of any inconsistency yields that each agent’s claims about herself must be the same as those of the agent to her left. That is, $p_{i,i+1}=p_{i+1,i+1}$ $\forall i$. Then, $\tau^{2}$ is inactive as well. Thus, there are no transfers.
4.6 Pure-Strategy Bayesian Nash Implementation
In this section, we take a brief digression to discuss pure-strategy implementation. While NPD was shown to be necessary for mixed-strategy implementation, a weaker version, called No Pure-Perfect Deceptions (NPPD) suffices for pure-strategy implementation. It is essentially identical to NPD except for the additional requirement that any deception $\alpha$ be degenerate, since playing a non-degenerate deception requires mixing by the agents. A pure-perfect deception is defined below.
Definition 9
A deception $\alpha=(\alpha_{i})_{i\in\mathcal{I}}$ is pure-perfect for state $s^{\prime}$ at state $s$ if it is a perfect deception for $s$ at $s^{\prime}$ and $\alpha$ is degenerate.
We now define the No Pure-perfect deceptions condition.
Definition 10
(Condition NPPD - No Pure-Perfect Deceptions) We say that a social choice function $f$ satisfies NPPD whenever for two states $s$ and $s^{\prime}$, if at $s$, state $s^{\prime}$ is nonrefutable, and there exists a pure-perfect deception for $s^{\prime}$, then $f(s)=f(s^{\prime})$.
We now state the pure-strategy implementation result.
Theorem 2
A social choice function is implementable in pure-strategy BNE if and only if it satisfies NPPD.
The proof of this result is derived as a simplification of our mixed-strategy implementation result, and is relegated to the online appendix.
No Pure-perfect deceptions is shown to be strictly weaker than NPD and strictly stronger than Stochastic Measurability in the online appendix. Therefore, in our setup, mixed-strategy implementation is a strictly harder problem than pure-strategy implementation. We note that this distinguishes this setup from the complete-information setup in Ben-Porath and Lipman (2012) where mixed-strategy Nash Implementation does not require a stronger condition on the social choice function. However, it is consistent with the rest of the Bayesian implementation literature though, in that mixed-strategy Nash implementation in Serrano and Vohra (2010) requires Mixed Bayesian Monotonicity, which is stronger than Bayesian Monotonicity which is required for pure-strategy implementation (Jackson (1991)).
5 Uncertain State Types
5.1 Setup
In this section, we study a more general model than the one considered above. Agents have a type $(s_{i},E_{i})$, among which the component $s_{i}\in S_{i}$ (hereafter the state type) represents the private information of the agent not associated with hard evidence - that is, the agent can arbitrarily misreport $s_{i}$ to any mechanism. $S_{i}$ is assumed to be finite. Denote by $s_{-i}$ the state type profile of the agents other than agent $i$, and let $S_{-i}=\Pi_{j\neq i}S_{j}$. The set of all states $S$ is given by $\Pi_{i\in I}S_{i}$, and a state $s$ is used to denote the state type profile $(s_{i})_{i\in\mathcal{I}}$.
The evidence realization $E_{i}$ is chosen by nature from the set of all possible evidence realizations for agent $i$, $\mathbb{E}_{i}$, and each $\mathbb{E}_{i}$ is assumed to be finite. Denote $\mathbb{E=}\Pi_{i\in\mathcal{I}}\mathbb{E}_{i}$. First, nature chooses a state $s=(s_{i})_{i\in\mathcal{I}}$ according to a prior $p^{s}\in\Delta S$.
We say that $s_{i}^{\prime}$ can be reached in $m$ steps from $s_{i}$ if there is a sequence $\left\{s_{j_{n}}^{n}\right\}_{n=1}^{m}$ such that $s_{j_{1}}^{1}=s_{i}$, $s_{j_{m}}^{m}=s_{i}^{\prime}$ and for each $n$, $j_{n}\in\mathcal{I}$, and for $n\geq 2$, agent $j_{n-1}$ of type $s_{j_{n-1}}$ believes that agent $j_{n}$ can be of type $s_{j_{n}}$ with positive probability. Define $\bar{S}_{i}(s)=\cup_{m=1}^{\infty}\left\{s_{i}^{\prime}:s_{i}^{\prime}\text{ can be reached in }m\text{ steps from }s_{i}\right\}$. Since the set of state types $S_{i}$ is assumed to be finite, there is a finite number of steps $k$ such that either it is not possible to reach a state type $s_{i}$ from another state type $s_{i}^{\prime}$, or it is possible to do so in less than $k$ steps. Let $\bar{S}(s)=\{\tilde{s}\in\Pi_{i\in\mathcal{I}}\bar{S}_{i}(s_{i}):p^{s}(\tilde{s})>0\}$. We say that a subset $\tilde{S}\subseteq S$ is belief-closed if for any $s\in\tilde{S}$, $\bar{S}(s)=\tilde{S}$.
After nature chooses $s$, evidence is realized according to a prior $p^{e}:S\rightarrow\Pi_{i\in\mathcal{I}}\Delta\mathbb{E}_{i}$. We denote by $p_{i}^{e}$ the evidence distribution of agent $i$. The domain of the SCF $f$ is given by $S=\Pi_{i\in\mathcal{I}}S_{i}$, so that $f:S\rightarrow A$. The critical difference between this and the previous setup is that the agents only know their own realization of $s_{i}$, but are unfamiliar with the others’ realizations, i.e. $s_{j}$. As before, conditional on the realized $s$, any agent’s evidence distribution is independent of the evidence distribution of any other agent.
5.2 Generalized NPD
[…]
We impose the information smallness assumption from McLean and Postlewaite (2002) in the following qualified sense - were the state types of agents other than agent $i$ to be $s_{-i}$, there is a small $\varepsilon_{I}>0$ and a unique state type $s_{i}$ of agent $i$ such that $p^{s}(s_{i},s_{-i})>1-\varepsilon_{I}$.  Further, given any $s_{i}$, there is a unique profile of opposing state types $s_{-i}$ such that agent $i$’s posterior belief about the state type profile $p_{s_{i}}^{s}$ satisfies $p_{s_{i}}^{s}(s_{i},s_{-i})>1-\varepsilon_{I}$. When these conditions are met, we say that each agent has an information size of $\varepsilon_{I}$. The results presented for the case where agents are informationally small will be in the context of large economies (formally characterized by $I\rightarrow\infty$), wherein we will further assume that $\varepsilon_{I}\rightarrow 0$ as $I\rightarrow\infty$. In the spirit of McLean and Postlewaite (2002), this is meant to reflect the notion that as the number of agents increases, each agent’s influence on the economy becomes progressively smaller. The main motivation behind this assumption is the case where there is negligible aggregate uncertainty, i.e. the state of nature can be inferred with high precision from all agents’ signals.
Note that in McLean and Postlewaite (2002), there is an “estimation” problem where conditionally i.i.d. signals received by agents are being used to estimate the state of the world. In their leading example for instance, as the number of agents increases, each agent contributes an increasingly smaller amount of information to the estimate of the state. An analogous interpretation in our setting is that the “signals” $s_{i}$ are being used to estimate the evidence distribution $p^{e}$, which directly corresponds to the state ($S$) of the world. In this view, the notion that each agent’s influence on the posterior estimate is small translates to the requirement that $\varepsilon_{I}\rightarrow 0$ as $I\rightarrow\infty$.
A strengthening of this assumption is nonexclusive information (Postlewaite and Schmeidler (1986), Vohra (1999)) which is obtained by setting $\varepsilon_{I}$ to 00 uniformly, so that aggregating the information from all but one agent identifies the state type of the remaining agent exactly. Note that with only two agents, nonexclusive information is identical to complete information, and with three or more agents, it is weaker than complete information.
In this setup, the notion of refutation is defined as follows - $E_{i}\downarrow s$ if $p_{i}^{e}(s)(E_{i}^{\prime})=0$ for any $E_{i}^{\prime}\supseteq E_{i}$. That is, whenever the state is $s$, agent $i$ is never endowed with $E_{i}$ or a superset thereof. [^17] Note that this leverages our assumption that $p^{e}$ yields uncorrelated evidence draws. This is because correlation among evidence draws might require agents to submit evidence profiles rather than evidence collections to refute a state. In the design of an implementing mechanism, this leads to a coordination issue since an agent may not have the incentive to submit their component of a refuting profile of evidence if others have not already submitted their components of this profile. Further, $s^{\prime}$ is nonrefutable at $s$ if any $E_{i}$ which some agent $i$ may be endowed with when the state is $s$ does not refute $s^{\prime}$.
As before, a deception for an agent $i$ is a function $\alpha_{i}:S_{i}\times\mathbb{E}_{i}\rightarrow\Delta(S_{i}\times\mathbb{E}_{i}\mathbb{)}$. We denote the state component of the deception (which is pure), i.e. $\alpha_{i}(s_{i},E_{i})(s_{i})$ by $\alpha_{i}^{s}:S_{i}\rightarrow S_{i}$, and the evidence component $\alpha_{i}(s_{i},E_{i})(E_{i})$ (which may involve mixing) by $\alpha_{i}^{e}:S_{i}\times\mathbb{E}_{i}\rightarrow\Delta\mathbb{E}_{i}$. [^18] The implementing mechanism we will create will prevent agents from mixing in the state dimension in any equilibrium. That is, the state component of the deception depends only on the state component of the type playing the deception. Deceptions must satisfy evidence feasibility, i.e. if $E_{i}^{\prime}\in\operatorname{supp}\alpha_{i}^{e}(s_{i},E_{i})$, then $E_{i}^{\prime}\subseteq E_{i}$. A deception can also be interpreted as a strategy in a direct revelation mechanism, which satisfies: (i) the evidence feasibility condition, i.e. when playing a deception, an agent cannot present articles of evidence she does not possess; and (ii) no mixing in the state dimension. Denote by $\sigma^{\ast}$ the truthtelling strategy in a direct mechanism. Now, we define a perfect deception.
Definition 11
(Perfect Deception) Let $\bar{S}(s)$ be the smallest belief-closed set containing $s$. A deception $\alpha$ is perfect for $s^{\prime}$ at $s$ if (i) for each agent $i$, $\alpha_{i}^{s}(s_{i})=s_{i}^{\prime}$; and (ii) for each state $\bar{s}\in\bar{S}(s)$ and for each type $\bar{s}_{i}\in\bar{S}_{i}(s)$ of each agent $i$, the belief induced under $\alpha$ on the messages of other agents in a direct mechanism is the same as that induced by $\sigma^{\ast}$ if her true state type were $\alpha_{i}^{s}(\bar{s}_{i})$, i.e. $\forall i,~{}\bar{s}_{i}\in\bar{S}_{i}(s)$,
$\sum_{\tilde{s}_{-i}\times\tilde{E}_{-i}}p^{s}(\bar{s}_{i},\tilde{s}_{-i})p_{-i}^{e}(\bar{s}_{i},\tilde{s}_{-i})(\tilde{E}_{-i})\alpha_{-i}(\tilde{s}_{-i},\tilde{E}_{-i})(\hat{s}_{-i},\hat{E}_{-i})$
$=$ $p^{s}(\alpha_{i}^{s}(\bar{s}_{i}),\hat{s}_{-i})p_{j}^{e}(\alpha_{i}^{s}(\bar{s}_{i}),\hat{s}_{-i})(\hat{E}_{-i})\text{ }\forall(\hat{s}_{-i},\hat{E}_{-i})\in S_{-i}\times\mathbb{E}_{-i}\text{.}$
On the RHS, $\sigma^{\ast}$ is used implicitly. Indeed, state type $\alpha_{i}^{s}(\bar{s}_{i},\sim)$ assumes that each opposing type $(\hat{s}_{-i},\hat{E}_{-i})$ reports their type truthfully, yielding the expression on the RHS.
We now spell out the key difference between the current setup and the base setup studied previously. In the base setup, given the true state $s^{\ast}$, every opposing type of other agents has the same belief over the evidence endowment of an agent $i$, whereas in the current setup, each opposing type may have different beliefs. For a deception to be perfect, it must adequately deceive each possible opposing type. This, in turn, requires us to consider the smallest belief-closed set containing the state $s$. We can now state the implementing condition.
Definition 12
(Condition GNPD - Generalized No Perfect Deceptions) SCF $f$ satisfies Generalized No Perfect Deceptions (GNPD) if $f(s)=f(s^{\prime})$ for any pair of states $s=(s_{i})_{i\in\mathcal{I}}$ and $s^{\prime}=(s_{i}^{\prime})_{i\in\mathcal{I}}$ such that (i) there is a perfect deception $\alpha$ for $s^{\prime}$ at $s$; and (ii) if $\bar{S}(s)$ is the smallest belief-closed set containing $s$, then $\alpha^{s}(\bar{s})$ is nonrefutable at $\bar{s}$ for any $\bar{s}\in\bar{S}(s)$.
Notice that to be perfect, a deception $\alpha$ must map the entire relevant subset of the state space (in this case $\bar{S}(s)$) to another subset. However, we single out a pair of states $s$ and $s^{\prime}$ so as to define the restriction on the SCF $f$.
The mechanism we use for establishing sufficiency of GNPD for implementation will extract up to $k$ orders of beliefs from each agent, and we remind the reader that $k$ is the minimum number of steps in which it is possible to reach any state type $s_{i}$ from any other state type $s_{i}^{\prime}$ unless it is not possible to do so in any number of steps. [^19] This is for simplicity. In practice, it may be possible to reduce the required level $k$ by focusing only on state type pairs which also satisfy the other requirements of the GNPD condition.
[…]
If a perfect deception does not exist at $s$ for $s^{\prime}$ then for any deception $\alpha$ there is a type $\tilde{s}_{i}$ of some agent $i$ which can be reached in $k$ or fewer steps from $s_{j}$ for some $j\in\mathcal{I}$ such that $\tilde{s}_{i}$ does not have the same beliefs under $\alpha$ as $\alpha_{i}^{s}(\tilde{s}_{i})$ would under truthtelling. In this case, we write $\tilde{s}_{i}\not\rightarrow\alpha_{i}^{s}(\tilde{s}_{i})$.
As in the model we studied earlier, the designer creates mechanisms $\mathcal{M}$ which are represented by tuples $(M,g,\tau)$, where $M$ is the message space, $g$ is the outcome function, and $\tau$ represents any transfers. The expected transfer in a mechanism is defined as follows.
Definition 13
Given a mechanism $\mathcal{M}=(M,g,\tau)$, the expected transfer $E[\tau]$ is given by $\sum_{s\in S}p^{s}(s)\sum_{E\in\mathbb{E}}p^{e}(s)(E)\sum_{m\in M}\sigma(s,E)(m)\tau(m)$.
That is, $E[\tau]$ is the ex-ante expected value of the (vector of) transfers. There are two possible notions of implementation used in this section, which we will define below.
Definition 14
A mechanism exactly implements an SCF $f$ in Bayesian Nash equilibrium if for any set of bounded utility functions, every equilibrium of the game induced by the mechanism yields the socially desirable outcome according to $f$ and yields zero transfers alongside.
Definition 15
[…]
Now, we state the main result of this section.
Theorem 3
Consider an economy with $I\geq 2$.
  • When agents are informationally small with information size $\varepsilon_{I}$, an SCF is implementable in a large economy (as $I\rightarrow\infty$) with vanishing expected transfers if it satisfies GNPD;
  • When information is nonexclusive, an SCF is exactly implementable in BNE with uncertain state types if and only if it satisfies GNPD.
The necessity proof and the mechanism used to establish sufficiency is presented in Appendix A.1, while the sufficiency proof is presented in the online appendix. The mechanism proposed in the online appendix is also shown to be fully continuous, which shows that it does not share the structure of integer or modulo games. In the following section, we provide a sketch of the proof alongside an illustrative figure to aid understanding.
5.3 Proof Sketch
Necessity follows from ideas similar to those in the base model. If a deception as required by the GNPD condition exists, then under this deception in state $s$, each type of each agent has the same beliefs and action set available to it as it does under truthtelling at $s^{\prime}$. Consequently, for any agent $i$ to play as if their state type were $s_{i}^{\prime}$ remains optimal, so that the outcome $f(s^{\prime})$ may arise in equilibrium.
For sufficiency, while following Theorem 1’s structure, uncertainty in state types presents novel challenges. First, a key issue emerges in the use of scoring rules and crosschecks in realigning towards the truth. In the base model, when an agent mixes, others expect inconsistency and present maximal evidence, enabling truthful predictions via scoring rules ($\tau^{2}$) and deterring mixing through crosschecks ($\tau^{3}$). However, with uncertain types, when type $s_{i}$ mixes, while direct opponents $s_{j}$ might present maximal evidence (expecting inconsistency due to $s_{i}$), their opponents $s_{k}$ may not, as they need not expect $s_{i}$. This breaks the chain of truthful predictions needed for effective crosschecks – a problem absent in the base model where mixing by “state type” $s^{\ast}$ guarantees all agents expect inconsistency with probability 1 since $s^{\ast}$ is common across agents.
Figure 2: A graphical depiction of the mechanism from the perspective of $s_{i}$.
In this extended model, we must ensure that all types in the belief chain ($s_{k}$ expected by $s_{j}$ expected by $s_{i}$) present maximal evidence, enabling truthful scoring rule reports from $s_{j}$ that ultimately discipline $s_{i}$’s outcome-determining report. [^21] A subtle detail is that $s_{j}$’s reports to the scoring rule may not be the exact truth, but rather some state type which induces the same distribution on opposing evidence as does the truth. Inferring the true profile of state types from the second report also requires $s_{i}$ to reveal her own type truthfully (in the distribution sense) in the second report. This occurs because when $s_{i}$ sets $c_{i}^{1}>0$, the types $s_{j}$ she expects also expect her and thus present maximal evidence, so that she predicts the maximal evidence of these opposing types. These details are formally outlined in Lemmas 4 and 5 in the online appendix. This is achieved by incorporating first and second order beliefs about inconsistency into the message space via the proper scoring rules $c^{1}$ and $c^{2}$ (which function somewhat like bets). [^22] For simplicity of illustration, we largely focus on the first two steps here, but the idea extends to third and higher order beliefs up to an order $k$, which is discussed above. Since $S$ is finite, $k$ is bounded. These “bets” and “bets-on-bets” about inconsistency enable evidence presentation incentivization at multiple belief orders. When $c^{1}$ or $c^{2}$ is positive, maximal evidence submission is guaranteed, helping realign reports toward truth as in the base mechanism. For clarity, this idea is also depicted graphically in Figure 2, from the perspective of type $s_{i}$ of agent $i$. The figure illustrates the steps by which the expectation of inconsistency, a refutation or a bet leads $s_{i}$ to reveal its state type truthfully.
Second, in this setting, perfect deceptions have an order component – a deception might work up to order $l$ but fail at the $k^{th}$ order of beliefs where some type’s true beliefs aren’t adequately mimicked. To ensure truthful realignment, we extract beliefs up to order $k$ using scoring rules $c^{3}$, $c^{4}$, …, $c^{k}$, conditioning evidence elicitation on positive $c^{k}$. When a type $s_{i}$ at the $k^{th}$ order isn’t perfectly deceived, it triggers a cascade: its bet causes types expecting it to set $c^{1}>0$, types (in turn) expecting these types to set $c^{2}>0$, and so on. This chain ensures maximal evidence submission from all types, enabling truthful realignment. This novel approach of ”carrying higher order beliefs down to the message space” addresses the challenge of multi-order deceptions.
Third, since agents do not know each other’s state types, we are faced with the question of how to “cancel out” the scoring rule transfer ($\tau^{2}$ in the base mechanism). Here, the assumptions on information (smallness or nonexclusiveness) allow us to (approximately, in the case of information smallness) infer the type of an agent from the reports by other agents, which allows for the cancellation even though an exact match might not be possible on the basis of only one other agent’s knowledge. [^23] In the case of information smallness, the cancellation happens with a large probability, which is why only the expected transfer is zero under information smallness. Unlike the base model though, this requires us to ensure that at least $I-1\,$agents are reporting their state types truthfully. We note here that the method to cancel out the scoring rules in equilibrium using a technique similar to that used in the proof of Theorem 1 is also used here to avoid transfers on equilibrium. This provides a solution to the issue described above.
6 Rationalizable Implementation in a general type space
In this section, we study a more general model than the one considered above. We allow agents to have a type $t_{i}\in T_{i}$ (where $T_{i}$ is finite) and allow agents’ types to be correlated. A type profile $t$ determines agents’ preferences over the outcomes in a set $A$, so that $\bar{u}_{i}(a,t)\,$ is agent $i$’s bounded utility when the outcome is $a$ and the type profile is given by $t$. Further, $u_{i}(a,\bar{\tau}_{i},t)=$ $\bar{u}_{i}(a,t)\,+\bar{\tau}_{i}$ is the quasilinear extension of $\bar{u}_{i}$. The type profile $t$ also determines agents’ evidence, which is represented by $\hat{E}_{i}(t_{i})\in\mathbb{E}_{i}$. States of the world are represented by type profiles $t=(t_{i})_{i\in\mathcal{I}}$. Each agent has a belief $q_{i}:T_{i}\rightarrow\Delta(T_{-i})$ where $T_{-i}=\Pi_{j\neq i}T_{j}$. Such a model is represented by $\mathcal{T=}(T_{i},\hat{E}_{i},q_{i})_{{}_{i\in\mathcal{I}}}$. A social choice function maps a state to the desirable outcome in that state, so that $f:T\rightarrow A$, where $T=\Pi_{i}T_{i}$. Once again, we are interested in implementing $f$ without relying on preference variation, so that we consider constant preferences as a possibility.
In this section, we will adapt the solution concept of Interim Correlated Rationalizability from Dekel, Fudenberg, and Morris (2007) to our setting. Consider a mechanism $\mathcal{M}=(M,g,\tau)$, which induces a static Bayesian game denoted by $G\left(\mathcal{M},\mathcal{T},u\right)$. Define the set of messages feasible for an agent $i$ of type $t_{i}$ by $M_{i,t_{i}}$. Denote by $R_{i}^{G}\left(t_{i}\right)$ the set of rationalizable messages for agent $i$ of type $t_{i}$. Then, we have $R_{i,0}^{G}\left(t_{i}\right)=M_{i,t_{i}}$;
$R_{i,k+1}^{G}\left(t_{i}\right)=\left\{\begin{array}[]{l|l}m_{i}\in M_{i,t_{i}}&\begin{array}[]{l}\text{ there exists }\pi_{i}\in\Delta\left(T_{-i}\times M_{-i}\right)\text{ such that }\\ (1)\pi_{i}\left(t_{-i},m_{-i}\right)>0\Rightarrow m_{-i}\in\Pi_{j\neq i}R_{j,k}^{G}\left(t_{j}\right)\\ (2)m_{i}\in\underset{m_{i}^{\prime}\in M_{i,t_{i}}}{\arg\max}\sum_{E_{-i},m_{-i}}u_{i}(g\left(m_{i}^{\prime},m_{-i}\right),\tau_{i}\left(m_{i}^{\prime},m_{-i}\right),t)\\ (3)\sum_{m_{-i}}\pi_{i}\left(t_{-i},m_{-i}\right)=q_{i}(t_{i})(t_{-i})\end{array}\end{array}\right\}\text{;}$
and $R_{i}^{G}\left(t_{i}\right)=\cap_{k=1}^{\infty}R_{i,k}^{G}\left(t_{i}\right)$. Then, denote by $R^{G}\left(t\right)=\Pi_{i\in I}R_{i}^{G}\left(t_{i}\right)$ the profile of all rationalizable messages.
With this solution concept, we formalize the implementation notion below.
Definition 16
A social choice function $f$ is rationalizably implementable with arbitrarily small transfers if for any $\varepsilon>0$, there is a mechanism $\mathcal{M}=(M,g,\tau)$ such that for any profile of bounded utility functions $\bar{u}=\left(\bar{u}_{i}\right)_{i\in\mathcal{I}}$, any type profile $t$, and any profile of rationalizable messages $m\in R^{G}\left(t\right)$ of the game $G\left(\mathcal{M},\mathcal{T},u\right)$, we have $g(m)=f(t)$. Further, $|\tau_{i}(m)|\leq\varepsilon$ for every message profile $m$.
That is, given an arbitrarily small (strictly) positive number $\varepsilon$, it should be possible to obtain the socially optimal outcome at any profile of rationalizable messages, and limit the transfers for any message profile (rationalizable or otherwise) to being no higher than $\varepsilon$. This represents a relaxation of the prior requirement that the transfers should be zero in equilibrium, but yields a relaxation of the solution concept to rationalizability, allows correlation among types, and (we will show) yields a weaker implementing condition than either NPD or NPPD.
We now describe the agent’s higher order beliefs. Since we wish to implement without relying on preference variation, we restrict agents to form beliefs only on evidence to obtain a necessary condition which is strong enough to allow implementation when preferences are constant. Note that the implementing mechanism we will present later will allow for preference variation, but will not rely on it for implementation. To this end, define $Z_{i}^{0}=\mathbb{E}_{-i}$, and for $k\geq 1$, define $Z_{i}^{k}=Z_{i}^{k-1}\times\Delta(Z_{-i}^{k-1})$. For each $i$, let $T_{i}^{\ast}$ denote the set of $i$’s collectively coherent hierarchies (Mertens and Zamir (1985)). For each $t_{i}\in T_{i}$, let $\hat{\pi}_{i}^{0}(t_{i})=\hat{E}_{i}(t_{i})$; construct mappings $\hat{\pi}_{i}^{k}:T_{i}\rightarrow\Delta(Z_{i}^{k-1})$ recursively for all $i\in\mathcal{I}$ and $k\geq 1$, such that $\hat{\pi}_{i}^{1}$ is the push-forward of $q_{i}(t_{i})$ given by the map from $T_{-i}$ to $Z_{-i}^{0}$, such that
$t_{-i}\longmapsto\hat{\pi}_{-i}^{0}(t_{-i})\text{,}$
and $\hat{\pi}_{i}^{k}(t_{i})$ is the push-forward of $q_{i}(t_{i})$ given by the map from $T_{-i}$ to $Z_{-i}^{k-1}$ such that
$t_{-i}\longmapsto(\hat{\pi}_{-i}^{0}(t_{-i}),\hat{\pi}_{-i}^{1}(t_{-i}),...,\hat{\pi}_{-i}^{k-1}(t_{-i}))\text{,}$
where for any $k=0,1,2,..$, $\hat{\pi}_{-i}^{k}(t_{-i})=\Pi_{j\neq i}\hat{\pi}_{j}^{k}(t_{j})$.
The mappings thus constructed, i.e. $\hat{\pi}_{i}^{\ast}(t_{i})=(\hat{\pi}_{i}^{0}(t_{i}),\hat{\pi}_{i}^{1}(t_{i}),...)$ assign to each type a hierarchy of beliefs. We now propose a new implementing condition which we term higher-order measurability (HOM).
Definition 17
A social choice function $f$ satisfies higher-order measurability if for any pair of states $t$ and $t^{\prime}$ such that $f(t)\neq f(t^{\prime})$, there is $k\in\mathbb{N}$ such that there is an agent $i$ whose $k^{th}$ order belief $\hat{\pi}_{i}^{k}(t_{i})\neq\hat{\pi}_{i}^{k}(t_{i}^{\prime})$.
The necessity of HOM is demonstrated in the online appendix. Intuitively, HOM requires that the social choice function $f$ be measurable on a partition of $T$ where each cell of the partition is such that if $t$ and $t^{\prime}$ belong to the same cell, then each agent has the same belief hierarchy in both $t$ and $t^{\prime}$.
We now consider the setup we work with for Bayesian Nash implementation. Suppose an SCF $f$ satisfies stochastic measurability, so that if $f(s)\neq f(s^{\prime})$, then there is an agent $i$ such that $p_{i}(s)\neq p_{i}(s^{\prime})$. Then, the first order belief of any agent $j\neq i$ is different between $s$ and $s^{\prime}$, so that $f$ must satisfy HOM. Conversely, if there are two states $s$ and $s^{\prime}$ so that for each agent $i$, $p_{i}(s)=p_{i}(s^{\prime})$, and evidence distributions are independent, then $f(s)$ must be identical to $f(s^{\prime})$ if $f$ is to satisfy HOM. Therefore, in such a setup, HOM is identical to stochastic measurability, and therefore weaker than either NPD or NPPD.
Towards establishing that HOM is also sufficient for implementation with the above implementation notion, we will define the relevant version of incentive compatibility in this setting. To do so, consider the direct mechanism with respect to $\mathcal{T=}(T_{i},\hat{E}_{i},q_{i})_{{}_{i\in\mathcal{I}}}$, where each agent $i$ makes a report $t_{i}\in T_{i}$ and the outcome chosen is $f((t_{i})_{i\in\mathcal{I}}$. No transfers are induced.
Definition 18
A social choice function $f$ satisfies evidence incentive compatibility if for each agent $i$, and each type $t_{i}\in T_{i}$,
$t_{i}\in\arg\max_{t_{i}^{\prime}\in\hat{T}_{i}(t_{i})}\sum_{t_{-i}\in T_{-i}}q_{i}(t_{i})(t_{-i})\bar{u}_{i}(f(t_{i}^{\prime},t_{-i}),(t_{i},t_{-i}))$
where $\hat{T}_{i}(t_{i})=\{t_{i}^{\prime}:\hat{E}_{i}(t_{i}^{\prime})\subseteq\hat{E}_{i}(t_{i})\}$.
That is, if agent $i$ is endowed with type $t_{i}$, and anticipates every other agent to report their type truthfully, then it is among her best responses to also report truthfully. This is consistent with Deneckere and Severinov (2008), who show that if agents can send every combination of their available messages, then the above condition is necessary and sufficient for truthtelling to be a Bayesian Nash equilibrium. Theorems 1 and 2, however, do not require this condition since the use of large transfers causes it to be automatically satisfied. Now, we present the main result of this section.
Theorem 4
A social choice function is rationalizably implementable with arbitrarily small transfers iff it satisfies higher-order measurability and evidence incentive compatibility.
The sufficiency of HOM is demonstrated through the construction of an implementing mechanism which is presented in the online appendix. Here, we describe the intuition behind the construction, which follows ideas from Abreu and Matsushima (1992). The first step is to use a small incentive to extract the evidence from each agent. Irrespective of the other reports they make, it will be a strictly dominant strategy to present all their evidence. In the second step, we ask each agent to predict the evidence presentation of the other agents and score their predictions using a quadratic scoring rule. The input to the scoring rule is the agent’s own type, since common knowledge of the type allows the designer to compute the posterior distribution $q_{i}(t_{i})$ and score the prediction appropriately. Since each agent expects every other agent will have presented all their evidence, she presents her true belief about the evidence presentations of the other agent in order to maximize the returns from this scoring rule. This elicits the other agent’s true first order belief. In the third step, we extract the agents true second order belief by asking them to predict the first order beliefs extracted in the previous step and scoring their predictions using a quadratic scoring rule. HOM assures us that if we repeat this process enough times, we will have extracted the true types of the agents. This mechanism performs the role of the dictatorial choice function in Abreu and Matsushima (1992). Each transfer here can be made small enough so that the overall transfer can be bounded below any preselected $\varepsilon$. In what follows, we refer to this extracted profile of types as $t^{0}$ and to agent $i$’s component of $t^{0}$ as $t_{i}^{0}$.
After achieving extraction, we proceed towards the goal of implementing the social choice function. The steps taken are essentially identical to those in Abreu and Matsushima (1992). Agents are asked for $J$ additional reports of their type, denoted $(t_{i}^{j})_{j=1,2,..,J}$ in what follows. Each profile of reports determines the outcome with a probability of $\frac{1}{J}$, but the following penalties are applied. To the first agent $i$ who deviates in their report $t_{i}^{j}$ from $t_{i}^{0}$, a penalty is applied which is large enough to dominate the incentive from affecting the outcome with a probability of $\frac{1}{J}$ but not so large as to incentivize lying in $t^{0}$. That such a balance can be achieved can be deduced from the fact that the designer can choose $J$ to be as large as required to reduce the incentive to affect the outcome. Then, no agent wants to be the first to deviate from the truthful report in $t^{0}$ and implementation obtains while the penalties remain switched off.
This mechanism however inherits some of the well known critiques of the Abreu-Matsushima type of mechanisms, most notably that it requires a high depth of reasoning, involves a fairly complex message space, and is more vulnerable to renegotiation (than say the mechanism proposed in the proof of Theorem 1) owing to the use of small transfers.
7 Related Literature
In this section, we discuss our paper’s results alongside some other related papers to provide the reader a comparison of our results against the literature. First, this paper is most closely related to the study of implementation with hard evidence, which was pioneered by Ben-Porath and Lipman (2012) and Kartik and Tercieux (2012) and later developed further in Banerjee, Chen, and Sun (2024b). When compared to this literature, our main contribution is twofold - the incorporation of uncertainty in the evidence dimension, and the incorporation of limited but nontrivial uncertainty in the payoff relevant state-types. [^24] Peralta (2019) also studies a variation of this problem but restricts attention only to direct mechanisms. We study the rich class of indirect mechanism which nevertheless do not depend on integer or modulo games.
In contrast to the deterministic evidence setting studied in Banerjee, Chen, and Sun (2024b), a measurability condition is insufficient for implementation due to the challenge posed by perfect deceptions. While under deterministic evidence, implementation only requires that some agent can separate states using evidence, with uncertain evidence the key factor becomes agents’ beliefs about evidence distributions rather than actual evidence endowments. This fundamental difference necessitates novel techniques including the construction of bet akin to Crémer and McLean (1988), scoring rules to extract agents’ beliefs about evidence distributions, and transfers that leverage on higher-order thinking. These techniques require careful design to avoid transfers occuring in equilibrium - challenges that do not arise in the deterministic case. [^25] In fact, due to these challenges, constructions in the style of Chen, Kunimoto, Sun, and Xiong (2022) which involve crosschecking agents using transfers instead of using integer or modulo games have not successfully been extended to cover incomplete information settings (even assuming nonexclusive information or information smallness).
Second, focusing on the classical bayesian implementation literature pioneered in Jackson (1991) and later developed in Serrano and Vohra (2010), our main contribution lies in extending these results to include evidence, allowing for implementation when preferences do not vary with the state. This allows the implementation literature to speak to important practical problems such as litigation, budget allocation etc. However, we acknowledge the following tradeoff - while previous articles on bayesian implementation allow for full bayesian uncertainty in state-types, we work with either nonexclusive information or informationally small agents. Further, the mechanisms in Jackson (1991) use modulo games and those in Serrano and Vohra (2010) use integer games. We present in this paper, to our knowledge, the first mechanism in the literature that achieves implementation in a finite mechanism in a Bayesian setting which explicitly accounts for mixed-strategy equilibria.
Third, compared to the literature on information smallness (e.g. McLean and Postlewaite (2002), McLean and Postlewaite (2004)), which study the conflict between incentive compatibility and efficiency, we study the implementation of general SCFs in such information environments. Worth noting is that while informationally small agents are compatible with incentive compatibility, the implementation of SCFs in such environments require that all but one agents be incentivized to tell the truth so that incentive compatibility can then be used to crosscheck the remaining agent. In the setting we study, this requires novel techniques such as bets and directly incorporating higher order beliefs into the message space - methodological innovations which provide mechanism designers with new tools. [^26] For instance, the latter finds use in Banerjee, Chen, and Sun (2024a), where we study implementation of SCFs in correlated equilibria.
More generally, Kartik and Tercieux (2012) provides an excellent survey of the mechanism design with evidence literature. An early reference in the field of mechanism design with evidence is Green and Laffont (1986) who study the principal-agent problem when the agent cannot manipulate the truth arbitrarily. This corresponds to a notion of evidence - an agent, by presenting messages which are only available to her in a certain set of states, can prove that the true state is within that set. Early contributions in the field of game theory and evidence include Milgrom (1981), Grossman (1981) and Dye (1985). More papers involving mechanism design and evidence include Bull and Watson (2007), Deneckere and Severinov (2008), Hart, Kremer, and Perry (2017), and Ben-Porath, Dekel, and Lipman (2019). Ben-Porath, Dekel, and Lipman (2021) study a setup involving stochastic evidence. While we assume that agents are endowed with evidence to begin with, Ben-Porath, Dekel, and Lipman (2021) study the process of evidence acquisition and the value of commitment. In addition, they focus on partial implementation whereas our focus is on full implementation.
Appendix A Appendix
A.1 Proof of Theorem 3
A.1.1 Proof of Necessity
Here we prove that any implementable SCF $f$ must satisfy GNPD. For contradiction suppose $f(s)\not=f(s^{\prime})$ for a pair of states $s$ and $s^{\prime}$ such that (i) there is a perfect deception $\alpha$ for $s^{\prime}$ at $s$; and (ii) if $\bar{S}(s)$ is the smallest belief-closed set containing $s$, then $\alpha^{s}(\bar{s})$ is nonrefutable at $\bar{s}$ for any $\bar{s}\in\bar{S}(s)$. Then, for each agent $i$, given type $\bar{s}_{i}\in\bar{S}_{i}(s)$, the belief induced by $\alpha$ on the messages of other agents in a direct mechanism is the same as that induced by $\sigma^{\ast}$ (the truthtelling strategy) if agent $i$ were of state type $\alpha_{i}^{s}(\bar{s}_{i})$. Denote the belief of such a type $\bar{s}_{i}$ by $b_{i}^{d}(\bar{s}_{i},\alpha)=(\sum_{\tilde{s}_{-i}\times\tilde{E}_{-i}}p^{s}(\bar{s}_{i},s_{-i})p_{-i}^{e}(\bar{s}_{i},s_{-i})(\tilde{E}_{-i})\alpha(\tilde{s}_{-i},\tilde{E}_{-i})(\hat{s}_{-i},\hat{E}_{-i}))_{(\hat{s}_{-i},\hat{E}_{-i})}$, where the superscript $d$ represents a direct mechanism. Consider any mechanism $\mathcal{M}=(M,g,\tau)$ which implements such an SCF, and consider any equilibrium $\sigma$ of the game induced by this mechanism.
Consider the strategies $\sigma\circ\alpha$ which denote the following strategy - each agent first plays the deception $\alpha$, and then plays the equilibrium strategy ($\sigma$) of the type that results from $\alpha$. We claim that this strategy is an equilibrium of the game but yields the outcome $f(s^{\prime})$ in state $s$.
To demonstrate this, we will first consider the incentives of an arbitrarily chosen type $(\bar{s}_{i},\bar{E}_{i})$ such that $\bar{s}_{i}\in\bar{S}_{i}(s)$. [^27] The set of types in $\bar{S}_{i}(s)$ are the only ones which are relevant for equilibrium analysis, since each agent expects types outside this set to appear with probability zero at any order of belief. Suppose $\alpha_{i}^{s}(\bar{s}_{i})=\hat{s}_{i}$. First, we note that the collection of evidence $\bar{E}_{i}$ (or a superset thereof) must also occur with positive probability when the state is given by $\hat{s}$, since otherwise this collection would refute $\hat{s}$ at $\bar{s}$. That is, the type $(\hat{s}_{i},\hat{E}_{i})$ for some $\hat{E}_{i}\supseteq\bar{E}_{i}$ occurs with positive probability when the state is $\hat{s}$. In this setting, this forms the analogue of condition (*) in the necessity proof of Theorem 1. The beliefs of the type $(\bar{s}_{i},\bar{E}_{i})$ over opposing messages in $\mathcal{M}$ under $\sigma\circ\alpha$ are given by
$b_{i}^{\mathcal{M}}(\bar{s}_{i},\sigma\circ\alpha)=\sum_{\tilde{s}_{-i}\times\tilde{E}_{-i}}p^{s}(\bar{s}_{i},\tilde{s}_{-i})p_{-i}^{e}(\bar{s}_{i},\tilde{s}_{-i})(\tilde{E}_{-i})\sigma(\alpha(\tilde{s}_{-i},\tilde{E}_{-i}))\text{.}$
[…]
Further, in this equilibrium, type $s_{i}\in\bar{S}_{i}(s)$ plays the equilibrium strategy of some type $(s_{i}^{\prime},\sim)$, so that the outcome in state $s$ is $f(s^{\prime})$ (since $\mathcal{M}$ implements $f$, and $\sigma$ is an equilibrium of the game induced by $\mathcal{M}$, $g(\sigma(s^{\prime}))=f(s^{\prime})$), which contradicts implementation.
A.1.2 Comparisons with Bayesian Monotonicity
Reinterpreted into our setting (with constant preferences), Bayesian monotonicity requires that for any deception $\alpha$ such that $f\left(\alpha(s)\right)\neq f(s)$ for some $s$, there is an agent $i$, a type $(s_{i},E_{i})$, and an alternate SCF (potentially including a transfer) $f^{\prime}$, such that $f^{\prime}\left(\alpha\left(s\right)\right)\succ_{i}f\left(\alpha\left(s\right)\right)$ while $f\left(s^{\prime}\right)\succsim_{i}f^{\prime}(\alpha_{i}(s_{i}^{\prime}),s_{-i}^{\prime})$. If no such SCF $f^{\prime}$ exists, then for $f$ to be implementable, it is necessary that $f\left(\alpha(s)\right)=f(s)$. We will show that if a deception $\alpha$ exists which is perfect for $s^{\prime}$ at $s$, it is not possible for such an $f^{\prime}$ to exist. Suppose for contradiction that it does. That is, there is a profitable deviation for type $(s_{i},E_{i})$ when the agents have played strategies which result in the outcome $f\left(\alpha\left(s\right)\right)=f(s^{\prime})$, i.e. $(s_{i},E_{i})$ can obtain an outcome she prefers to $f\left(\alpha\left(s\right)\right)=f(s^{\prime})$. Since $s^{\prime}$ is nonrefutable at $s$, $E_{i}$ or a superset exists when the profile of state types is $s^{\prime}$. Then, this type can also profitably deviate from the desired outcome $f\left(s^{\prime}\right)$, which contradicts $f\left(s^{\prime}\right)\succsim_{i}f^{\prime}(\alpha_{i}(s_{i}^{\prime}),s_{-i}^{\prime})$.
A.1.3 Sufficiency Mechanism
A Challenge Scheme
The mechanism involves agents placing bets to signal that a deception has been played through a challenge scheme. First, denote by $\mathcal{A}_{i}$ the set of all possible deceptions $\alpha_{i}$ for some agent $i$, and let the set $\mathcal{A}=\Pi_{i\in\mathcal{I}}\mathcal{A}_{i}$ denote the set of all possible deceptions. Fix a state type $s_{i}$ of some agent $i$. Given $s_{i}$, let $p_{\alpha_{-i},s_{i}}(s_{-i},E_{-i})$ denote the probability that agent $i$ of state type $s_{i}$ ascribes to other agents (denoted $-i$) presenting a message profile $(s_{-i},E_{-i})$ in a direct mechanism if they are playing the profile of deceptions $\alpha_{-i}$. Mathematically,
$p_{\alpha_{-i},s_{i}}^{e}(s_{-i},E_{-i})=\sum_{s_{-i}}p^{s}(s_{i},s_{-i})\sum_{E_{-i}^{\prime}\in\mathbb{E}_{-i}}p_{i}^{e}(s_{i},s_{-i}^{\prime})(E_{-i}^{\prime})\alpha_{-i}(s_{-i}^{\prime},E_{-i}^{\prime})(s_{-i},E_{-i})\text{.}$
Let $p_{\alpha_{-i},s_{i}}=(p_{\alpha_{-i},s_{i}}^{e}(s_{-i},E_{-i}))_{{}_{(s_{-i},E_{-i})\in S_{-i}\times\mathbb{E}_{-i}}}$ be the probabilities that agent $i$ of state type $s_{i}$ ascribes to each message profile of agents $-i$ when agents $-i$ play the deception $\alpha_{-i}$. $P_{s_{i}}=(p_{\alpha_{-i},s_{i}})_{\alpha_{-i}\in\mathcal{A}_{-i}}$ represents the set of all possible vectors of probabilities as described above for any deception the other agents can play. Finally, let $p_{s_{i}}^{\ast}$ be the vector of probabilities that agent $i$ of state type $s_{i}$ ascribes to each message profile of agents $-i$ when agents $-i$ play truthfully.
We remind the reader that we write $s_{i}\not\rightarrow s_{i}^{\prime}$ when there is no deception $\alpha$ such that the type $s_{i}$ has the same beliefs under $\alpha$ as she would if agents were playing according to the truthtelling strategy $\sigma^{\ast}$, and her true type were $s_{i}^{\prime}$. The challenge scheme is presented in the following lemma.
Lemma 3
For any agent $i$, there is a finite set $B_{i}=\{b_{i}(s_{i},s_{i}^{\prime}):s_{i}\not\rightarrow s_{i}^{\prime}\}$ such that $b_{i}(s_{i},s_{i}^{\prime})\cdot p_{s_{i}}^{\ast}<0$ and $b_{i}(s_{i},s_{i}^{\prime})\cdot p>0~{}\forall p\in P_{s_{i}}$.
Proof. Since $s_{i}\not\rightarrow s_{i}^{\prime}$, $p_{s_{i}^{\prime}}^{\ast}\notin P_{s_{i}}$. Further, $P_{s_{i}}$ is closed and convex. Then, from the separating hyperplane theorem, there is a hyperplane $b_{i}$ such that $b_{i}\cdot p_{s_{i}^{\prime}}^{\ast}<0$ and $b_{i}\cdot p>0$ $\forall p\in P_{s_{i}}$.
For a pair of state types $s_{i}$ and $s_{i}^{\prime}$ one such $b_{i}$ is sufficient. Iterating over all possible pairs of state types $s_{i},s_{i}^{\prime}$ for which which $s\not\rightarrow s^{\prime}$ yields the set $B_{i}$. Therefore, each $B_{i}$ is finite.
Here we say that an agent $i$ of type $s_{i}$ cannot be “deceived” by the others in terms of her belief about the message profiles of other agents. Since state types permit arbitrary misreporting, if such a perfect deception is not possible, it must be because the other agents are not possessed of the required deception in the evidence dimension.
Message Space
The message space is given by $M_{i}=S_{i}\times S_{i}\times E_{i}\times[0,1]\times...\times[0,1]\times S_{i}$, with a typical message being represented by $m_{i}=(s_{i}^{1},s_{i}^{2},E_{i},c_{i}^{1},c_{i}^{2},...,c_{i}^{k},c_{i}^{k+1},c_{i}^{k+2},s_{i}^{b})$. [^28] While this choice of message space yields an infinite mechanism, the methodology of the proof is not similar to the use of integer or modulo games. Indeed, this is for expositional simplicity, and a technique similar to that used in the treatment of costly evidence in Banerjee, Chen, and Sun (2024b) can be used to discretize the values of $c_{i}^{1}$ and $c_{i}^{2}$. The reader is reminded that $k$ is the minimum number of steps in which it is possible to reach any state type $s_{i}$ from any other state type $s_{i}^{\prime}$ unless it is not possible to do so at all. The profile of messages $(m_{i})_{i\in\mathcal{I}}$ is denoted by $m$.
Outcome
The outcome is given by $f((s_{i}^{1})_{i\in\mathcal{I}})$ when the profile of state types $s^{1}$ is such that $p^{s}(s^{1})>1-\varepsilon_{I}$. [^29] The reader is reminded that $\varepsilon_{I}>0$ is the information size of agents. Otherwise, it is a preselected arbitrary outcome $a\in A$.
Transfers
The transfers involved in this mechanism are described below. First, we define some preliminaries. We derive a profile $\bar{s}$ from $s^{2}=(s_{i}^{2})_{i\in\mathcal{I}}$ as follows: if $p^{s}(s^{2})=0$, choose an arbitrary $\bar{s}\in S$ such that $p_{j,\bar{s}_{i}}^{e}=p_{j,s_{i}^{2}}^{e}$ $\forall j$ if it is feasible, otherwise let $\bar{s}=s^{2}$. If , $p^{s}(s^{2})>0$, let $\bar{s}=s^{2}$.
A message profile $m$ is said to be inconsistent if $p^{s}((s_{i}^{1})_{i\in\mathcal{I}})<1-\varepsilon_{I}$ OR $s^{1}\neq\bar{s}$, otherwise it is consistent. Given a profile of state types $s_{-i}$ which is consistent with the prior $p^{s}$, $\tilde{s}_{i}(s_{-i})$ represents the unique $s_{i}$ such that $p^{s}(s_{i},s_{-i})>1-\varepsilon_{I}$. [^30] Informational smallness with information size $\varepsilon_{I}$ guarantees the existence of such $\tilde{s}_{i}$.
A message profile $m$ is said to involve a refutation if some $E_{i}\downarrow s^{1}$. A message profile $m$ is said to involve a bet if $s_{i}^{b}\neq s_{i}^{1}$ for some $i$. Further, given a message profile $m$, we say that only agent $i$ has placed a bet if $s_{i}^{b}\neq s_{i}^{1}$ while $s_{j}^{b}=s_{j}^{1}$ for each $j\neq i$. We say that an agent $i$ of type $s_{i}$ expects inconsistency, refutation or a bet on some message $m_{i}$ if given the strategy of the agents, and her beliefs about the types of the other agents, she expects with positive probability a profile of opposing messages $m_{-i}$ such that $m=(m_{i},m_{-i})$ is inconsistent, involves a refutation or a bet respectively.
As in the base mechanism, the first transfer provides an agent the incentive to submit evidence when there is inconsistency, or an active bet, or if a profile has been refuted by an agent (or if these issues are expected at an order of belief up to $k$). However, the incentive structure requires some precise design which is formally defined below:
[…]
where $\delta>0$ is a small positive number, and $\zeta_{-i}(c)=\Pi_{j\neq i}\Pi_{\hat{k}=1}^{k+2}c_{j}^{\hat{k}}$, i.e. the product of every other agent’s $c$ values. [^31] The mechanism also implements with $\zeta_{-i}(c)$, but we add this component to make the mechanism fully continuous, a point which is discussed later.
When active, $\tau^{1}$ provides an incentive for the agent to provide more evidence. If any of the conditions in the first part of the definition are met, i.e. if agent $i$ has refuted $s^{1}$ (when $s^{1}$ is consistent) or $m$ involves inconsistency, then for agent $i$, $\tau^{1}$ provides this incentive.
The second transfer extracts the first order beliefs of agents regarding inconsistency, bets or refutations. First, define a random variable $R_{i}^{1}$as follows.
Ri1(m)={
1, if m is inconsistent, involves a refutation by another agent ji,
or a bet placed by another agent ji
0 otherwise.
[…]
We remind the reader that we use the notation $Q(p,x)=[2p(x)-p\cdot p]$ to score a prediction $p$ against a realization $x$ of a random variable $X$. When $X$ is a Bernoulli random variable, i.e. when it takes values in the set $\{0,1\}$, the scoring rule can be normalized such that the prediction-realization pair $(0,0)$ (which stands for the prediction that the probability of $X=1$ is zero with the realization $X=0$) is scored as 00 in the following way - $\tilde{Q}(p,x)=Q(p,x)-Q(0,0)$, where with some abuse of notation, $p$ is the predicted probability of the realization $x$. Then, we define $\tau^{2}$ as follows:
$\tau_{i}^{2,1}\left(m\right)=\frac{\delta}{I}\tilde{Q}(c_{i}^{1},R^{1}(m))\text{.}$
Note that in equilibrium, this yields no transfer since each $c_{i}^{1}=0$ (agents do not expect $R_{i}^{1}=1$), and $R_{i}^{1}=0$.
The next transfer extracts agents’ beliefs about $c_{i}^{k-1}$, or (equivalently) their $k^{th}$ order beliefs about inconsistency, active bets (by others), or refutations (also by others). Define a random variable $R^{k-1}$ which is set to $1$ if the message profile $m$ involves any $c_{.}^{k-1}>0$ and 00 otherwise. Then,
$\tau_{i}^{2,k}\left(m\right)=\frac{\delta}{I}\tilde{Q}(c_{i}^{k},R^{k-1}(m))\text{.}$
Once again, this transfer is designed to be off in equilibrium.
The next transfer asks agents to predict the evidence presentation of other agents using a scoring rule, and has a term appended to ensure no equilibrium transfers. $Q(s_{i},E_{-i})$ gives the value of the scoring rule when other agents have presented the evidence profile $E_{-i}$ and agent $i$’s prediction is $p_{s_{i}}^{e\ast}$. [^32] The reader is reminded that $p_{s_{i}}^{e\ast}$ is the vector of probabilities that agent $i$ of state type $s_{i}$ ascribes to each evidence profile $E_{-i}$ of agents $-i$ when agents $-i$ reveal their type truthfully. With $\tau\geq 1$, define
$\tau_{i}^{3}(m)=\frac{\delta}{I}[Q(s_{i}^{2},E_{-i})-Q(\tilde{s}_{i}(s_{-i}^{1}),E_{-i})]\text{.}$
The next transfer is a cross check between the other agents’ second reports and agent $i$’s first report. Recall that $\bar{s}$ represents a consistent version of any report profile $s^{2}$ if such a profile is possible. Define
[…]
Note that for each $\bar{s}$, there exists a unique $\tilde{s}_{i}(\bar{s}_{-i})$ under information smallness with size $\varepsilon_{I}$.
Finally, the following transfer allocates the gains or losses from any bets which are placed.
τi5(m)={
δIbi(si1,sib)(s-i,E-i), if sibsi1 and ↛sibsi1
0 otherwise.
[…]
We remind the reader that the bet $b_{i}(s_{i}^{1},s_{i}^{b})$ is defined in Lemma 3.
Transfers $\tau^{1}$, $\tau^{2,.}$, $\tau^{3}$ and $\tau^{5}$ are at a scale of $\frac{\delta}{I}$, so that a transfer of $1$ dollar can dominate both the outcome and any losses from these transfers.
We draw the reader’s attention to two critical points - (i) an agent’s evidence submission only affects their utility through the transfer $\tau^{1}$; and (ii) the agent’s choice of $c^{1}$, $c^{2}$ etc. only affects their utility through $\tau^{2}$ since $\tau^{1}$ only depends on the $c^{.}$ submissions of other agents.
Appendix B Evidence, Refutation and Deceptions
Consider any two states $s$ and $s^{\prime}$. Given an evidence structure, the following table captures the possibilities regarding refutations and deceptions, with the final column describing whether the states are separable in terms of a mechanism or not. When states are not separable by a mechanism, an SCF must prescribe the same social choice outcomes if it is implementable. With some abuse of notation, $s\downarrow s^{\prime}$ denotes that $s$ is refutable at $s^{\prime}$ and $s\rightarrow s^{\prime}$ denotes that there is a perfect deception from $s$ to $s^{\prime}$.
Case $s\downarrow s^{\prime}$ $s^{\prime}\downarrow s$ $s\rightarrow s^{\prime}$ $s^{\prime}\rightarrow s$ Description
1 0 0 0 0 No perfect deceptions: states are separable
2 0 0 0 1 $s^{\prime}$ nonrefutable at $s$ and $s^{\prime}\rightarrow s$: not separable
3 0 0 1 0 $s$ nonrefutable at $s^{\prime}$ and $s\rightarrow s^{\prime}$: not separable
4 0 0 1 1 $s\rightarrow s^{\prime}$ and $s^{\prime}\rightarrow s$ and no refutations: not separable
5 0 1 0 0 No perfect deceptions: states are separable
6 0 1 0 1 Impossible, see below
7 0 1 1 0 $s$ nonrefutable at $s^{\prime}$ and $s\rightarrow s^{\prime}$: not separable
8 0 1 1 1 Impossible.
9 1 0 0 0 No perfect deceptions: states are separable
10 1 0 0 1 $s^{\prime}$ nonrefutable at $s$ and $s^{\prime}\rightarrow s$: not separable
11 1 0 1 0 Impossible
12 1 0 1 1 Impossible
13 1 1 0 0 Mutual refutation: states are separable
14 1 1 0 1 Mutual refutation: Impossible
15 1 1 1 0 Mutual refutation: Impossible
16 1 1 1 1 Mutual refutation: Impossible
Cases 6, 8, 11 and 12: If $s^{\prime}$ is refutable at $s$ but $s$ is not refutable at $s^{\prime}$, there are articles of evidence available at $s$ which are not available at $s^{\prime}$, so that a perfect deception cannot exist from $s^{\prime}$ to $s$.
Cases 14, 15, and 16 are impossible because under a mutual refutation, there are articles of evidence in either state which do not exist in the other state so that a perfect deception cannot exist.
Appendix C Connection to Ben-Porath and Lipman (2012)
Ben-Porath and Lipman (2012) use a notion of nomenclature to associate articles of evidence with subsets of the state space. In essence, an article of evidence is associated with the set of states in which it occurs. A natural generalization to the setting we study here would be to associate an article of evidence with the set of states in which it occurs with positive probability. When probability distributions are degenerate (so that we have a deterministic evidence setting), it is clear that only one collection may occur with positive (and indeed unit) probability. We denote this collection by $E_{i}(s)$. With this notion of nomenclature, we state the analogous conditions from Ben-Porath and Lipman (2012) below.
  1. 1.
    (e1)$\ \forall e_{i}\in E_{i}(s)$, $s\in e_{i}$
  2. 2.
    (e2) $e_{i}\in E_{i}(s)\implies e_{i}\in E_{i}(s^{\prime})$ for every $s^{\prime}\in e_{i}$.
The following proposition proves that (se1) and (se2) are indeed generalizations of (e1) and (e2).
Proposition 1
If $p$ is degenerate, then (se1) and (se2) are identical to (e1) and (e2) respectively.
Proof. First, consider some $e_{i}$ such that $s\notin e_{i}$. Then, any collection $E_{i}$ containing $e_{i}$ refutes $s$. From se1 therefore, $p_{i}(s)(E_{i})=0$ for any $E_{i}\ni e_{i}$. Then, $E_{i}\neq E_{i}(s)$, whereupon we have obtained that if $s\notin e_{i}$, then $e_{i}\notin E_{i}(s)$, which is (e1). Thus, (se1) implies (e1).
Next, suppose $E_{i}\downarrow s$. Then, $\exists e_{i}\in E_{i}$ such that $s\notin e_{i}$. From (e1), $e_{i}\notin E_{i}(s)$, so that $p_{i}(s)(E_{i})=0$, which establishes (se1). Therefore, (e1) implies (se1) as well.
Now, we turn to (se2). Consider a pair of states $s$ and a collection $E_{i}$ such that $E_{i}\not\downarrow s^{\prime}$. Notice that under degenerate probabilities, $p_{i}(s^{\prime})(E_{i}^{\prime})>0\implies E_{i}^{\prime}=E_{i}(s^{\prime})$ and $p_{i}(s)(E_{i})>0\implies E_{i}=E_{i}(s)$, so that (se2) translates to the following: If $\forall e_{i}\in E_{i}(s)$, if $s^{\prime}\in e_{i}$, then $E_{i}(s^{\prime})\supseteq E_{i}(s)$. Now, consider any $e_{i}\in E_{i}(s)$ such that $s^{\prime}\in e_{i}$. Then, $E_{i}(s^{\prime})\supseteq E_{i}(s)\implies e_{i}\in E_{i}(s^{\prime})$, which establishes (e2).
Finally, suppose $E_{i}\not\downarrow s^{\prime}$ and $p_{i}(s)(E_{i})>0$. Under degenerate probabilities, this yields $E_{i}(s)\not\downarrow s^{\prime}$, so that $\forall e_{i}\in E_{i}(s)$, $s^{\prime}\in e_{i}$. From (e2), $e_{i}\in E_{i}(s^{\prime})$. Since this is true for all $e_{i}$, we have $E_{i}(s^{\prime})\supseteq E_{i}(s)$, which is identical to (se2) under degenerate probabilities. We have, therefore established that (se1) and (se2) correspond to (e1) and (e2) under degenerate probabilities.
Next, we consider the following question: Since articles can be named according to the subset of states in which they occur with positive probability, does it suffice to allow an agent to reveal a “most informative” article of evidence rather than presenting an entire collection? More precisely, consider the following example evidence structure [^33] Please note that this example varies from the running example in its structure. It is an abstract example intended for illustrative purposes.:
State/Agent A B
$U$$\{\{L,M,H,U\}\}$$\{\{L,M,H,U\}\}$
$H$ $0.3\times\{\{L,M,H,U\},\{M,H\}\}+0.7\times\{\{L,M,H\},\{M,H\}\}$$\{\{L,M,H,U\}\}$
$M$ $0.5\times\{\{L,M,H,U\},\{M,H\}\}+0.5\times\{\{L,M,H\},\{M,H\}\}$$\{\{L,M,H,U\}\}$
$L$$\{\{L,M,H\},\{L,M,H,U\}\}$$\{\{L,M,H,U\}\}$
In both states $M$ and $H$, agent A, the only agent with evidence variation always has the article $\{M,H\}$ and this is the most informative article which is available to her. If we were to only allow her to present one article of evidence, then there would be a perfect deception in both directions between the states $M$ and $H$ so that we would not be able to distinguish them using any mechanism. However, if we were to allow the submission of entire collections of evidence, then there is no perfect deception from state $M$ to $H$ (the article $\{L,M,H\}$ is not available often enough) and no perfect deception from state $H$ to $M$ (the article $\{L,M,H,U\}$ is not available often enough). Therefore, we can distinguish between these states. Thus, there is a loss of generality when restricting attention to the most informative article of evidence. In a deterministic evidence setting though, this restriction is without loss of generality, as can be seen in both Ben-Porath and Lipman (2012) and Banerjee, Chen, and Sun (2024b).
Appendix D NPPD vs NPD and Stochastic Measurability
In this section, we demonstrate that NPPD is strictly stronger than Stochastic Measurability and strictly weaker than NPD. To see that NPPD is strictly stronger than Stochastic Measurability, consider the example below.
State/Agent A B
$U$ $0.1\times\{\{L,M,H,U\}\}$$\{\{L,M,H,U\}\}$
$+0.9\times\{\{L,M,H,U\},\{M,H,U\},\{H,U\}\}$
$H$ $0.1\times\{\{L,M,H,U\},\{M,H,U\}\}$$\{\{L,M,H,U\}\}$
$+0.9\times\{\{L,M,H,U\},\{M,H,U\},\{H,U\}\}$
$M$ $0.5\times\{\{L,M,H,U\},\{M,H,U\}\}+0.5\times\{\{L,M,H,U\}\}$$\{\{L,M,H,U\}\}$
$L$$\{\{L,M,H,U\}\}$$\{\{L,M,H,U\}\}$
Clearly, $U$ is nonrefutable at $H$. Consider the deception at state $H$ given by
$\alpha_{A}(H,\{\{L,M,H,U\},\{M,H,U\}\})=(U,\{\{L,M,H,U\}\})\text{;}$
$\alpha_{A}(H,\{\{L,M,H,U\},\{M,H,U\},\{H,U\}\})=(U,\{\{L,M,H,U\},\{M,H,U\},\{H,U\}\}).$
While the SCF satisfies stochastic measurability, this is a pure-perfect deception.
To see that NPPD is strictly weaker than NPD, we refer back to the example in Section 2 of the main paper. We reproduce the evidence structure below:
State Evidence (Firm A)
$H$ $0.6\times\{\{M,H\},\{L,M,H\}\}+0.4\times\{\{L,M,H\}\}$
$M$ $0.4\times\{\{M,H\},\{L,M,H\}\}+0.6\times\{\{L,M,H\}\}$
$L$$\{\{L,M,H\}\}$
It is easy to see that there is no pure-perfect deception in the above example - at $H$, the type $\{\{M,H\},\{L,M,H\}\}$ must present all their evidence, as type $\{\{L,M,H\}\}$ does not have enough evidence to mimic $\{\{M,H\},\{L,M,H\}\}$ at $M$, but then we induce too much probability (0.6 instead of 0.4) on $\{\{M,H\},\{L,M,H\}\}$. This demonstrates that NPPD is strictly weaker than NPD.
Appendix E Proof of Theorem 2
The necessity of NPPD stems from arguments similar to those for the proof of necessity of NPD for mixed implementation. Sufficiency is established by referring to the mechanism used in the proof of Theorem 1. With the exception of the bets, the rest of the mechanism remains identical. Note that Claims 1 and 2 make no use of either the condition NPD or nontrivial randomization by agents. Further, Claims 1 and 2 imply that there is no equilibrium where agents present refutable lies. This leaves us with profiles consistent with nonrefutable lies.
So suppose agents present a profile consistent with a nonrefutable lie $s^{\prime}$. In contrast to the arguments in Section 4.2, where it was sufficient for an agent to inform the designer of the true state, in this paradigm, depending on the strategy profile, it is possible for different evidence collections to be low or high in probability relative to the expected probabilities in state $s^{\prime}$. Therefore, it is impossible for the designer to place an appropriate bet if she is informed only of the true state. Since $S$ and $\mathbb{E}$ are finite, the number of pure deceptions is finite, since any pure deception profile is a map $\alpha:S\times\mathbb{E}\rightarrow$ $S\times\mathbb{E}$. Therefore, they can be identified with numbers from $1$ through $Z$ where $Z$ is the total number of possible pure deceptions.
We augment the message space of the base mechanism by adding the strategy identifier $z\in(1,2,..,Z)$. Once informed of the deception identifier, the designer can identify an agent $i$ and a pair of collections $E_{i}^{\prime}$ and $E_{i}^{\prime\prime}$ such that $p_{\alpha_{i}(s)}(E_{i}^{\prime})<p_{i}(s^{\prime})(E_{i}^{\prime})$ and $p_{\alpha_{i}(s)}(E_{i}^{\prime\prime})>p_{i}(s^{\prime})(E_{i}^{\prime\prime})$ where $\alpha_{i}$ is the pure deception being played by agent $i$. Then, there are numbers $\gamma<0$ and $\delta>0$ such that (i) $\gamma\times p_{\alpha_{i}(s)}(E_{i}^{\prime})+\delta\times p_{\alpha_{i}(s)}(E_{i}^{\prime\prime})>0$ and (ii) $\gamma\times p(s^{\prime})(E_{i}^{\prime})+\delta\times p(s^{\prime})(E_{i}^{\prime\prime})>0$. These numbers $\gamma$ and $\delta$ form the bets in this setting. It is profitable to place a bet because of inequality (i). The rest of the analysis remains the same as earlier, in that the bet induces evidence submission and triggers a realignment through Lemma 2 in the main text. Therefore, agents must present truthful reports and we obtain implementation. Note that there are no bets in equilibrium as betting against the truth yields losses (from inequality (ii)).
Appendix F Proof of Theorem 3
F.1 Proof of Sufficiency
The proof utilizes the following strategy - first, we prove the part of Theorem 3 related to information smallness, and then present the modifications to the proof which yield the result under nonexclusive information.
F.1.1 A Challenge Scheme
As before, the mechanism will involve agents placing bets to signal that a deception has been played which can be detected. This requires the construction of a challenge scheme. First, denote by $\mathcal{A}_{i}$ the set of all possible deceptions $\alpha_{i}$ for some agent $i$, and let the set $\mathcal{A}=\Pi_{i\in\mathcal{I}}\mathcal{A}_{i}$ denote the set of all possible deceptions. Fix a state type $s_{i}$ of some agent $i$. Given $s_{i}$, let $p_{\alpha_{-i},s_{i}}(s_{-i},E_{-i})$ denote the probability that agent $i$ of state type $s_{i}$ ascribes to other agents (denoted $-i$) presenting a message profile $(s_{-i},E_{-i})$ in a direct mechanism if they are playing the profile of deceptions $\alpha_{-i}$. Mathematically,
$p_{\alpha_{-i},s_{i}}^{e}(s_{-i},E_{-i})=\sum_{s_{-i}}p^{s}(s_{i},s_{-i})\sum_{E_{-i}^{\prime}\in\mathbb{E}_{-i}}p_{i}^{e}(s_{i},s_{-i}^{\prime})(E_{-i}^{\prime})\alpha_{-i}(s_{-i}^{\prime},E_{-i}^{\prime})(s_{-i},E_{-i})\text{.}$
Let $p_{\alpha_{-i},s_{i}}=(p_{\alpha_{-i},s_{i}}^{e}(s_{-i},E_{-i}))_{{}_{(s_{-i},E_{-i})\in S_{-i}\times\mathbb{E}_{-i}}}$ represent the vector of probabilities that agent $i$ of state type $s_{i}$ ascribes to each message profile of agents $-i$ when agents $-i$ play the deception $\alpha_{-i}$. Further, let $P_{s_{i}}=(p_{\alpha_{-i},s_{i}})_{\alpha_{-i}\in\mathcal{A}_{-i}}$, the set of all possible vectors of probabilities of the type described above for any deception the other agents can play. Finally, let $p_{s_{i}}^{\ast}$ be the vector of probabilities that agent $i$ of state type $s_{i}$ ascribes to each message profile of agents $-i$ when agents $-i$ reveal their type truthfully.
We remind the reader that we write $s_{i}\not\rightarrow s_{i}^{\prime}$ when there is no deception $\alpha$ such that the type $s_{i}$ has the same beliefs under $\alpha$ as she would if agents were playing according to the truthtelling strategy $\sigma^{\ast}$, and her true type were $s_{i}^{\prime}$. With these preliminaries established, we now present the challenge scheme in the following lemma.
Lemma 3
For any agent $i$, there is a finite set $B_{i}=\{b_{i}(s_{i},s_{i}^{\prime}):s_{i}\not\rightarrow s_{i}^{\prime}\}$ such that $b_{i}(s_{i},s_{i}^{\prime})\cdot p_{s_{i}}^{\ast}<0$ and $b_{i}(s_{i},s_{i}^{\prime})\cdot p>0~{}\forall p\in P_{s_{i}}$.
Proof. Since $s_{i}\not\rightarrow s_{i}^{\prime}$, $p_{s_{i}^{\prime}}^{\ast}\notin P_{s_{i}}$. Further, $P_{s_{i}}$ is closed and convex. Then, from the separating hyperplane theorem, there is a hyperplane $b_{i}$ such that $b_{i}\cdot p_{s_{i}^{\prime}}^{\ast}<0$ and $b_{i}\cdot p>0$ $\forall p\in P_{s_{i}}$.
For a pair of state types $s_{i}$ and $s_{i}^{\prime}$ one such $b_{i}$ is sufficient. Iterating over all possible pairs of state types $s_{i},s_{i}^{\prime}$ for which which $s\not\rightarrow s^{\prime}$ yields the set $B_{i}$. Therefore, each $B_{i}$ is finite.
Note that here we say that an agent $i$ of type $s_{i}$ cannot be “deceived” by the others in terms of her belief about the message profiles of other agents. However, state types permit arbitrary misreporting (lying). Then, if such a perfect deception is not possible, it must be because the other agents are not possessed of the required deception in the evidence dimension.
F.2 Mechanism
Now, we construct a mechanism to implement social choice functions which satisfy GNPD. First, we define the message space of the mechanism.
F.2.1 Message Space
The message space of the mechanism is given by $M_{i}=S_{i}\times S_{i}\times E_{i}\times[0,1]\times...\times[0,1]\times S_{i}$, with a typical message being represented by $m_{i}=(s_{i}^{1},s_{i}^{2},E_{i},c_{i}^{1},c_{i}^{2},...,c_{i}^{k},c_{i}^{k+1},c_{i}^{k+2},s_{i}^{b})$. [^34] While this choice of message space yields an infinite mechanism, the methodology of the proof is not similar to the use of integer or modulo games. Indeed, this is for expositional simplicity, and a technique similar to that used in the treatment of costly evidence in Banerjee, Chen, and Sun (2024b) can be used to discretize the values of $c_{i}^{1}$ and $c_{i}^{2}$. The reader is reminded that $k$ is the minimum number of steps in which it is possible to reach any state type $s_{i}$ from any other state type $s_{i}^{\prime}$ unless it is not possible to do so in any number of steps. The profile of messages $(m_{i})_{i\in\mathcal{I}}$ is denoted by $m$.
F.2.2 Outcome
The outcome is given by $f((s_{i}^{1})_{i\in\mathcal{I}})$ when the profile of state types $s^{1}$ is such that $p^{s}(s^{1})>1-\varepsilon_{I}$. [^35] The reader is reminded that $\varepsilon_{I}>0$ is the information size of agents. Otherwise, it is a preselected arbitrary outcome $a\in A$.
F.2.3 Transfers
The transfers involved in this mechanism are described below. First, we define some preliminaries. We derive a profile $\bar{s}$ from $s^{2}=(s_{i}^{2})_{i\in\mathcal{I}}$ as follows: if $p^{s}(s^{2})=0$, choose an arbitrary $\bar{s}\in S$ such that $p_{j,\bar{s}_{i}}^{e}=p_{j,s_{i}^{2}}^{e}$ $\forall j$ if it is feasible, otherwise let $\bar{s}=s^{2}$. If , $p^{s}(s^{2})>0$, let $\bar{s}=s^{2}$.
A message profile $m$ is said to be inconsistent if $p^{s}((s_{i}^{1})_{i\in\mathcal{I}})<1-\varepsilon_{I}$ OR $s^{1}\neq\bar{s}$, otherwise it is consistent. Given a profile of state types $s_{-i}$ which is consistent with the prior $p^{s}$, $\tilde{s}_{i}(s_{-i})$ represents the unique $s_{i}$ such that $p^{s}(s_{i},s_{-i})>1-\varepsilon_{I}$. [^36] Informational smallness with information size $\varepsilon_{I}$ guarantees the existence of such $\tilde{s}_{i}$.
A message profile $m$ is said to involve a refutation if some $E_{i}\downarrow s^{1}$. A message profile $m$ is said to involve a bet if $s_{i}^{b}\neq s_{i}^{1}$ for some $i$. Further, given a message profile $m$, we say that only agent $i$ has placed a bet if $s_{i}^{b}\neq s_{i}^{1}$ while $s_{j}^{b}=s_{j}^{1}$ for each $j\neq i$. We say that an agent $i$ of type $s_{i}$ expects inconsistency, refutation or a bet on some message $m_{i}$ if given the strategy of the agents, and her beliefs about the types of the other agents, she expects with positive probability a profile of opposing messages $m_{-i}$ such that $m=(m_{i},m_{-i})$ is inconsistent, involves a refutation or a bet respectively. We now define the transfers.
As in the base mechanism, the first transfer provides an agent the incentive to submit evidence when there are issues with the submitted profile of messages - i.e. there is inconsistency, or an active bet, or if a profile has been refuted by an agent (or if these issues are expected at an order of belief up to $k$). However, the incentive structure requires some precise design which is formally defined below:
[…]
where $\delta>0$ is a small positive number, and $\zeta_{-i}(c)=\Pi_{j\neq i}\Pi_{\hat{k}=1}^{k+2}c_{j}^{\hat{k}}$, i.e. the product of every other agent’s $c$ values. [^37] The mechanism also implements with $\hat{c}=1$, but we add this component to make the mechanism fully continuous, a point which is discussed later.
In any of the parts, when active, $\tau^{1}$ provides an incentive for the agent to provide more evidence. If any of the conditions in the first part of the definition are met, i.e. if agent $i$ has refuted $s^{1}$ (when $s^{1}$ is consistent) or $m$ involves inconsistency, then for agent $i$, $\tau^{1}$ provides a positive incentive. Similarly, if agent $i$ is the only agent who has placed a bet, then the incentive is positive as well. These conditions take precedence over the negative incentive, which is active only when the previous conditions have not been met, and $\tau^{1}$ is active for agent $i$ because other agents have set $c_{j}^{1}>0$ or $c_{j}^{2,l}>0$ for any $l$. Notice that agent $i$ has no control over $\tau^{1}$ through $c^{1}$ and $c^{2,.}$, rather, when $\tau^{1}$ is active through these channels, the agent can only minimize her losses by presenting all the evidence she is endowed with.
The second transfer extracts the first order beliefs of agents regarding inconsistency, bets or refutations. First, define a random variable $R_{i}^{1}$as follows.
[…]
We remind the reader that we use the notation $Q(p,x)=[2p(x)-p\cdot p]$ to score a prediction $p$ against a realization $x$ of a random variable $X$. When $X$ is a Bernoulli random variable, i.e. when it takes values in the set $\{0,1\}$, the scoring rule can be normalized such that the prediction-realization pair $(0,0)$ (which stands for the prediction that the probability of $X=1$ is zero with the realization $X=0$) is scored as 00 in the following way - $\tilde{Q}(p,x)=Q(p,x)-Q(0,0)$, where with some abuse of notation, $p$ is the predicted probability of the realization $x$. Then, we define $\tau^{2}$ as follows:
$\tau_{i}^{2,1}\left(m\right)=\frac{\delta}{I}\tilde{Q}(c_{i}^{1},R^{1}(m))\text{.}$
Note that in equilibrium, this yields no transfer since each $c_{i}^{1}=0$ (agents do not expect $R_{i}^{1}=1$), and $R_{i}^{1}=0$.
The next transfer extracts agents’ beliefs about $c_{i}^{k-1}$, or (equivalently) their $k^{th}$ order beliefs about inconsistency, active bets (by others), or refutations (also by others). Define a random variable $R^{k-1}$ which is set to $1$ if the message profile $m$ involves any $c_{.}^{k-1}>0$ and 00 otherwise. Then,
$\tau_{i}^{2,k}\left(m\right)=\frac{\delta}{I}\tilde{Q}(c_{i}^{k},R^{k-1}(m))\text{.}$
Once again, this transfer is designed to be off in equilibrium.
The next transfer asks agents to predict the evidence presentation of other agents using a scoring rule, and has a term appended to ensure no equilibrium transfers. $Q(s_{i},E_{-i})$ gives the value of the scoring rule when other agents have presented the evidence profile $E_{-i}$ and agent $i$’s prediction is $p_{s_{i}}^{e\ast}$. [^38] The reader is reminded that $p_{s_{i}}^{e\ast}$ is the vector of probabilities that agent $i$ of state type $s_{i}$ ascribes to each evidence profile $E_{-i}$ of agents $-i$ when agents $-i$ reveal their type truthfully. With $\tau\geq 1$, define
$\tau_{i}^{3}(m)=\frac{\delta}{I}[Q(s_{i}^{2},E_{-i})-Q(\tilde{s}_{i}(s_{-i}^{1}),E_{-i})]\text{.}$
The next transfer is a cross check between the other agents’ second reports and agent $i$’s first report. Recall that $\bar{s}$ represents a consistent version of any report profile $s^{2}$ if such a profile is possible. Define
τi4(m)={
−1, if si1~si(¯s-i)
0 otherwise.
[…]
Note that for each $\bar{s}$, there exists a unique $\tilde{s}_{i}(\bar{s}_{-i})$ under information smallness with size $\varepsilon_{I}$.
Finally, the following transfer allocates the gains or losses from any bets which are placed.
[…]
We remind the reader that the bet $b_{i}(s_{i}^{1},s_{i}^{b})$ is defined in Lemma 3.
Transfers $\tau^{1}$, $\tau^{2,.}$, $\tau^{3}$ and $\tau^{5}$ are at a scale of $\frac{\delta}{I}$, so that a transfer of $1$ dollar can dominate both the outcome and any losses from these transfers.
Before the main proof, we draw the reader’s attention to two critical points - (i) an agent’s evidence submission only affects their utility through the transfer $\tau^{1}$; and (ii) the agent’s choice of $c^{1}$, $c^{2}$ etc. only affects their utility through $\tau^{2}$ since $\tau^{1}$ only depends on the $c^{.}$ submissions of other agents.
F.3 Proof of Continuity
Claim 5
Under constant preferences, the proposed mechanism yields fully continuous utilities for the agents under the discrete metric on the state type space and evidence space and the usual metric on $c_{i}^{k}$ and transfers.
Proof. When the preferences are constant over the space of outcomes, agent’s utilities are fully determined by the transfers. Consider the message space for agent $i$ which is given by $(s_{i}^{1},s_{i}^{2},E_{i},c_{i}^{1},c_{i}^{2},...,c_{i}^{k},c_{i}^{k+1},c_{i}^{k+2},s_{i}^{b})$. Components other than $c_{i}^{.}$ are either state type or evidence reports on which any function is trivially continuous under the discrete metric. It remains to show that the transfers are continuous in $c_{i}^{.}$. The message components $c_{i}^{.}$ affect agent utilities only through $\tau^{1}$ and $\tau^{2}$. The second transfer $\tau^{2}$ is a quadratic scoring rule, which is continuous in its inputs. Further, an agent $i$, when making a small change to any $c_{i}^{.}$ cannot affect her own utility through $\tau^{1}$. She can only affect the utilities of the others. Consider an agent $j\neq i$ such that $\tau_{j}^{1}=\frac{\delta\hat{c}_{i}}{I}\times|E_{i}|$ where $\hat{c}=\Pi_{l\neq i}\Pi_{\hat{k}=1}^{k+2}c_{l}^{k}$. Since $\tau^{1}$ is a product of the $c$ values of the other agents multiplied with a constant, it is clear that this function too is continuous in $c_{i}^{.}$.
F.4 Proof of Implementation
Fix an arbitrary profile of state types given by $s=(s_{i})_{i\in\mathcal{I}}$. The proof explicitly involves belief hierarchies, so that it is convenient to coin the phrase +expects to denote the phrase “expects with positive probability”. For instance,“every opposing state type $s_{j}$ that type $s_{i}$ believes exists with positive probability” may alternately be written as “every $s_{j}$ which $s_{i}$ +expects”. On the other hand, the word “expects” is used to denote “believes with probability 1”. First we prove two lemmas which find use in multiple places in the proof.
Lemma 4
If an agent $i$ of state type $s_{i}$ expects inconsistency, refutation or an active bet in some message $m_{i}$ on the support of her mixed strategy, then (i) each type $s_{k}$ which is +expected by any $s_{j}$ which (in turn) is +expected by $s_{i}$ presents maximal evidence, (ii) $s_{i}$ sets either $c_{i}^{1}>0$ or $c_{i}^{2}>0$ in $m_{i}$.
Proof. If agent $i$ of state type $s_{i}$ expects inconsistency or a refutation by another agent or a bet by another agent on a message $m_{i}$, it is optimal for her to set $c_{i}^{1}>0$ on $m_{i}$ (owing to $\tau^{2,1}$). Every type $s_{j}$ of every other agent $j$ which she +expects in turn +expect her and thus set $c_{j}^{2}>0$ (for optimality under $\tau^{2,2}$). Further, these types $s_{j}$ also present maximal evidence since they +expect $s_{i}$ and thus expect $c_{i}^{1}>0$. Further, agent $i$ expects that any types $s_{k}$ of agents $k\neq j$ which these $s_{j}$ +expect in turn expect $c_{j}^{2}>0$ and thus provide maximal evidence (owing to $\tau^{1}$).
If instead agent $i$ of type $s_{i}$ performs a refutation or places a bet using $s_{i}^{b}$, then each type $s_{j}$ which $s_{i}$ +expects in turn +expects the refutation or the bet and hence sets $c_{j}^{1}>0$ for optimality under $\tau^{2}$. Each type $s_{k}$ which is +expected by these types $s_{j}$ in turn expect $c_{j}^{1}>0$ and thus present maximal evidence for optimality under $\tau^{1}$. Further, since the types $s_{j}$ which are +expected by $s_{i}$ set $c_{j}^{1}>0$, $s_{i}$ sets $c_{i}^{2}>0$ for optimality under $\tau^{2}$.
Figure 3: An illustration of Lemmas 4 and 5.
Lemma 5
If an agent $i$ of state type $s_{i}$ (i) expects each type $s_{k}$ which is +expected by any $s_{j}$ which (in turn) is +expected by $s_{i}$ to present maximal evidence and (ii) sets $c_{i}^{1}>0$ or $c_{i}^{2}>0$ in some of her messages, then type $s_{i}$ presents the truth in $s_{i}^{1}$ in every message.
Proof. Since agent $i$ of type $s_{i}$ expects every $s_{j}$ which she +expects to (in turn) expect that all opposing evidence $E_{-j}$ is maximal, she expects that $s_{j}^{2}$ must be such that $p_{k,s_{j}^{2}}^{e}=p_{k,s_{j}}^{e}$ $\forall k\neq j$ for optimality under $\tau^{3}$. Further, type $s_{i}$ also chooses $s_{i}^{2}$ such that $p_{j,s_{i}^{2}}^{e}=p_{j,s_{i}}^{e}$ $\forall j\neq i$ since each opposing type she +expects also +expects $s_{i}$ and thus expects that either $c_{i}^{1}>0$ or $c_{i}^{1}>0$ (by hypothesis) and thus presents maximal evidence. Then, from information smallness, type $s_{i}$ expects with probability greater than $1-\varepsilon_{I}$ that there is a unique choice of $\bar{s}$ available, and for this $\bar{s}$, $\tilde{s}_{i}(\bar{s}_{-i})=s_{i}$. [^39] Recall that while other profiles may happen with small proabbility, consitency is defined as having a probability of $1-\varepsilon_{I}$ or more. Therefore, it is a profitable deviation to to switch any untrue $s_{i}^{1}$ to $s_{i}$, i.e. the truth.
Remark 3
An alternate view of Lemmas 4 and 5 is that Lemma 4 tells us that for any type $s_{i}$, the expectation of inconsistency, refutation or an active bet has two consequences - (i) it incentivizes her to set $c_{i}^{1}>0$, and (ii) it causes her to expect that types twice removed from her in her hierarchy of beliefs present maximal evidence. Lemma 5 tells us that these two consequences in turn have the effect of inducing agent $i$ to report truthfully in $s_{i}^{1}$. Visually, Figure 4 shows these interactions as an input output map.
Figure 4: Input Output Map: Lemmas 4 and 5.
Claim 6
In any equilibrium, no agent mixes in $s^{1}$.
Proof. Suppose for contradiction that agent $i$ of type $s_{i}$ mixes in $s_{i}^{1}$. From information smallness, there is a unique profile of opposing state types $s_{-i}$ which type $s_{i}$ expects with a probability at least $1-\varepsilon_{I}$. Then, she expects that at most one of her state type reports is consistent with the opposing messages so that type $s_{i}$ expects inconsistency in some of her messages. Lemmas 4 and 5 in turn yield that she sets $s_{i}^{1}=s_{i}$ in every message, so that it is not possible that she mixes, which is a contradiction.
Claim 7
In every profile of messages $m$ on the support of any equilibrium, $s^{1}$ is such that $p^{s}(s^{1})>0$.
Proof. Suppose not, that is, there is a message profile $m$ on the support of some equilibrium such that $p^{s}(s^{1})=0$. Suppose that this message profile occurs with positive probability when the profile of state types is given by $s=(s_{i})_{i\in\mathcal{I}}$. Consider an arbitrary agent $i$ of type $s_{i}$. This type $s_{i}$ +expects the profile $s$, and thus expects inconsistency in some of her messages (since $01-\varepsilon_{I}>$ 0). [^40] Note that $p^{s}(s^{1})=0$ implies inconsistency under our definition. Therefore she sets $s_{i}^{1}=s_{i}$ in every message (Lemmas 4 and 5). Since $i$ and $s_{i}$ are arbitrarily chosen, each agent presents the truth in $s^{1}$, which contradicts $p^{s}(s^{1})=0$, since $p^{s}(s)>0$.
Claim 8
In every profile of messages $m$ on the support of any equilibrium, $s^{1}$ is not a refutable lie.
Proof. Suppose for contradiction that there is a message profile $m$ on the support of some equilibrium in which $s^{1}$ a refutable lie, and that this message profile $m$ occurs with positive probability when the profile of state types is given by $s=(s_{i})_{i\in\mathcal{I}}$. Without loss of generality, suppose type $s_{i}$ may have an article of evidence $E_{i}$ with positive probability such that $E_{i}\downarrow s^{1}$. Optimality under $\tau^{1}$ implies that $s_{i}$ presents this evidence when endowed with it. Then, for each agent $j\neq i$, the type $s_{j}$ expects a refutation, sets $c_{j}^{1}>0$, and from Lemmas 4 and 5 presents the truth.
Further, for optimality under $\tau^{2}$, $s_{i}$ sets $c_{i}^{2}>0$. This causes each type $s_{j}^{\prime}$ for each agent $j\neq i$ which is +expected by $s_{i}$ to set $c_{j}^{3}>0$ which in turn causes each type $s_{k}$ which is (in turn) +expected by these $s_{j}$ to present maximal evidence. This satisfies the antecedent of Lemma 5 for type $s_{i}$, so that it may only present the truth in $s^{1}$. In combination with the fact that each $s_{j}$ must present the truth, this yields a contradiction - if $s_{i}$ and each $s_{j}$ for $j\neq i$ presents the truth in $s^{1}$, then $s^{1}$ cannot be a refutable lie.
Claim 9
In every profile of messages $m$ on the support of any equilibrium, in state $s=(s_{i})_{i\in\mathcal{I}}$, $s^{1}$ is not a nonrefutable lie such that $f(s^{1})\neq f(s)$.
Proof. Suppose to the contrary that when the true profile of state types is $s$, $s^{1}$ is a nonrefutable lie such that $f(s^{1})\neq f(s)$ for some profile of messages $m$ in the support of an equilibrium. From the fact that $f$ satisfies GNPD, we have that there is a type $\tilde{s}_{i}$ of some agent $i$ which can be reached in $k$ or fewer steps starting from some $s_{j}$ ($j\in\mathcal{I}$) such that the belief of this type on the opposing message components $(s_{-i}^{1},E_{-i})$ under the strategies being played is different from its true beliefs $p_{\tilde{s}_{i}}^{\ast}$. [^41] We remind the reader that $k$ is the minimum number of steps in which it is possible to reach any state type $s_{i}$ from any other state type $s_{i}^{\prime}$ unless it is not possible to do so in any number of steps. It is optimal for this agent to place a bet by setting $s_{i}^{b}=\tilde{s}_{i}$ since $s_{i}^{b}$ influences type $\tilde{s}_{i}$’s utility only through $\tau^{1}$ and $\tau^{5}$ and $\tau^{1}$ yields a reward from betting, and $\tau^{5}$ also yields an expected reward. Since $\tilde{s}_{i}$ appears in the $k^{\text{th}}$ order belief of $s_{j}$, $c_{j}^{k}>0$. Then, each type $s_{l}$ that $s_{j}$ +expects presents maximal evidence and sets $c_{l}^{k+1}$ and each type that these $s_{l}$ in turn +expect present maximal evidence, so that each such $s_{l}$ presents messages in $s^{2}$ which satisfy $p_{n,s_{l}^{2}}^{e}=p_{n,s_{l}}^{e}$ $\forall n\neq l$. Further, since each $s_{l}$ presents maximal evidence, $s_{j}$ presents messages in $s^{2}$ which satisfy $p_{l,s_{j}^{2}}^{e}=p_{l,s_{j}}^{e}$ $\forall l\neq j$. Together, this implies that for optimality under $\tau^{4}$, $s_{j}^{1}$ is the truth.
Every other $s_{i}$ for $i\neq j$ sets $c_{i}^{k+1}>0$ since $c_{j}^{k}>0$, and each $s_{j}^{\prime}$ which is +expected by $s_{i}$ presents maximal evidence and sets $c_{j}^{k+2}>0$, so that each each $s_{l}$ which is, in turn +expected by these $s_{j}^{\prime}$ present maximal evidence. This implies that each $s_{j}^{\prime}$ presents messages in $s^{2}$ which satisfy $p_{l,s_{j}^{2}}^{e}=p_{l,s_{j}^{\prime}}^{e}$ $\forall l\neq j$. Further, $s_{i}$ also presents messages in $s^{2}$ which satisfy $p_{j,s_{i}^{2}}^{e}=p_{j,s_{i}}^{e}$ $\forall j\neq i$. Together, this implies that for optimality under $\tau^{4}$, $s_{i}^{1}$ is the truth.
However, this means that each $s_{i}$ in the state-type profile $s$ presents the truth, so that $s$ cannot be a nonrefutable lie.
Claim 10
The mechanism yields the outcome $f(s)$, and the expected transfers are vanishing in equilibrium in a large economy.
Proof. From the previous claims, in any state $s$, $s^{1}$ is such that $f(s^{1})=f(s)$, since $s^{1}$ is consistent (Claim 7) but not with either a refutable or a non-refutable lie which induces a different outcome (Claims 8 and 9). It remains to show that the expected transfer is zero in any equilibrium when $I\rightarrow\infty$. It is easy to see that each transfer other than $\tau^{4}$ vanishes when $I\rightarrow\infty$. [^42] In each case, the numerator is a finite number while the denominator is $I$.
To show that $\tau^{4}$ yields a zero expected transfer, we will first show that for any realization of type profiles, $\tau^{4}$ is active with a probability of at most $\varepsilon_{I}$. If $\tau^{4}$ is active on a message profile $m$ which occurs on the support of an equilibrium in some state $s$, then it must be the case that for some type $s_{i}$ of some agent $i$, $s_{i}^{1}\neq\tilde{s}_{i}(\bar{s}_{-i})$, i.e. there is inconsistency. Consider an arbitrary type $s_{j}$ of some arbitrary agent $j$. Type $s_{j}$ +expects the profile $s$ and thus +expects inconsistency and from Lemma 4, presents maximal evidence and sets $c_{j}^{1}>0$ or $c_{j}^{2}>0$. Further, $s_{j}$ expects all opposing types to +expect her, and thus +expect $c_{j}^{1}>0$ or $c_{j}^{2}>0$, and thus present maximal evidence. Type $s_{j}$ therefore sets $s_{j}^{2}$ such that $p_{k,s_{j}^{2}}^{e}=p_{k,s_{j}}^{e}$ $\forall k\neq j$ for optimality under $\tau^{3}$. But, $j$ and $s_{j}$ are arbitrarily chosen and if all types of all agents report a type in their second report which induces the same belief on the opposing evidence as does the truth, then $\tau_{i}^{4}$ can be active (nonzero) only if the realized type profile $s$ is such that for some agent $i$, $s_{i}\neq\tilde{s}_{i}(s_{-i})$. Consider a partition of $S$ into $S_{1}$ and $S_{2}$ such that $S_{2}=\{s\in S:\exists i$ s.t. $s_{i}\neq\tilde{s}_{i}(s_{-i})\}$, and $S_{1}=S\backslash S_{2}$. The expected transfer from $\tau^{4}$ can then be written as
$E[\tau^{4}]=p^{s}(S_{1})\times 0+\sum_{s\in S_{2}}p^{s}(s)\sum_{E\in\mathbb{E}}p^{e}(s)(E)\sum_{m\in M}\sigma(s,E)(m)\tau^{4}(m)\text{.}$
Note that $\sum_{s\in S_{2}}p^{s}(s)<\varepsilon_{I}$ since agents are informationally small with information size $\varepsilon_{I}$ and $\varepsilon_{I}\rightarrow 0$ as $I\rightarrow\infty$ in a large economy. Since $\tau^{4}(m)$ is finite for each $m$, $\lim_{I\rightarrow\infty}E[\tau^{4}]=0$, and thus $\lim_{I\rightarrow\infty}E[\tau]=0$.
Claim 11
When information is nonexclusive, the mechanism yields the outcome $f(s)$, and there are no transfers in equilibrium.
Proof. Under the stronger assumption of nonexclusive information, Lemmas 4 and 5, and Claims 6 through  9 remain true. This yields that $s^{1}$ is such that $f(s^{1})=f(s)$ so that the mechanism yields the outcome $f(s)$ when the profile of state types is $s$. It remains to establish that there are no transfers. First, we establish that there can be no inconsistency. [^43] Recall that a message profile $m$ is said to be inconsistent if $p^{s}((s_{i}^{1})_{i\in\mathcal{I}})=0$ OR $s^{1}\neq s^{2}$. If $s_{i}^{1}\neq\tilde{s}_{i}(\bar{s}_{-i})$ for some $i$ on a message profile $m$ in the support of some equilibrium, then since type $s_{i}$ expects the profile $s$, and hence expects inconsistency, by Lemma 4, she expects every type $s_{j}$ of other agents $j$ which she +expects to present maximal evidence, so that (by $\tau^{3}$) $s_{i}^{2}$ is such that $p_{j,s_{i}^{2}}^{e}=p_{j,s_{i}}^{e}$ $\forall j\neq i$. Consider any other agent $j$ and the type $s_{j}$. This type +expects the profile $s$ and therefore +expects inconsistency. By Lemma 4, $s_{j}$ expects every type of other agents which she +expects to present maximal evidence, so that $s_{j}^{2}$ is such that $p_{k,s_{j}^{2}}^{e}=p_{k,s_{j}}^{e}$ $\forall k\neq j$. This yields that there is a unique $\bar{s}$ ($=s$) which is consistent with $s^{2}$, so that $s_{i}^{1}=\tilde{s}_{i}(\bar{s}_{-i})$, a contradiction. Thus, agents are consistent in any message $m$. This yields that $\tau^{3}$ and $\tau^{4}$ are inactive in equilibrium. If, in state $s$, any agent $i$ places a bet in equilibrium, other agents expect the bet with positive probability (since they +expect the profile $s$) and present maximal evidence. Therefore, it is not optimal to place a bet, so that $\tau^{5}$ is inactive in equilibrium. Further, since no refutations or bets are possible, agents have a strict incentive to report $c_{i}^{1}=c_{i}^{2}=...=c_{i}^{k}=0$. Therefore, $\tau^{1}$, $\tau^{2,.}$, and $\tau^{3}$ are also inactive in equilibrium.
Appendix G Proof of Theorem 4
The reader will recall that higher-order measurability requires that the social choice function $f$ be measurable on a partition of $T$ where each cell of the partition is such that if $t$ and $t^{\prime}$ belong to the same cell, then each agent has the same belief hierarchy in both $t$ and $t^{\prime}$. We will show that higher-order measurability is necessary for rationalizable implementation with arbitrarily small transfers regardless of the mechanism in use. Accordingly, fix a mechanism $\mathcal{M}=(M,g,\tau)$.
A general type space model (see for instance P2008) is defined as $\mathcal{T}=(T_{i},\hat{\theta}_{i},q_{i})_{i\in\mathcal{I}}$. The function $\hat{\theta}_{i}$ maps each type to a payoff type. A payoff type consists of any constituent component of an agent’s type which affects how she evaluates an allocation. In general, such models assume that all the actions are available to the agent in each state. In our case though, evidence is involved, which may vary in its availability with the state. Recalling that we work with utilities which are bounded by $1$ dollar, we associate the following cost structure with evidence in order to model evidence availability. Suppose the mechanism $\mathcal{M}$ induces a maximum transfer of $\bar{U}$. If an agent $i$ of type $t_{i}$ presents only such articles as are available to her in order to obtain an allocation $a$ and transfer $\tau_{i}$, then her utility is given by $u_{i}(a,\tau_{i},t)$. If an agent “presents” an unavailable article of evidence to do the same, her utility is given by $u_{i}(a,\tau_{i},t)-2\bar{U}-1$. Since it is never worthwhile to obtain an outcome if it involves a $12\bar{U}-$ 1 dollar cost in order to do so, an agent never does so, so that we have effectively modelled evidence availability. Further, we avoid depending on preference variation for implementation, so that $u_{i}(a,\tau_{i},t)=u_{i}(a,\tau_{i},t^{\prime})$ $\forall t,t^{\prime}\in T$, i.e. preferences are constant. With these two restrictions then, only evidence is payoff relevant, so that in our case, $\hat{\theta}_{i}:T_{i}\rightarrow\mathbb{E}_{i}$, so that $\hat{\theta}_{i}(t_{i})=\hat{E}_{i}(t_{i})$. Then, a general type space model yields our model $\mathcal{T=(}T_{i},\hat{E}_{i},q_{i})_{{}_{i\in\mathcal{I}}}$.
That higher-order measurability is necessary for implementation in rationalizable strategies in then immediate from Proposition 2 in P2008 - if the set of rationalizable strategies is identical for each type of each agent, then different outcomes cannot be achieved in different states. Alternately, any profile of strategies that is rationalizable at $t$ is also rationalizable at $t^{\prime}$ if $\hat{\pi}_{i}^{k}(t_{i})=\hat{\pi}_{i}^{k}(t_{i}^{\prime})$ for all $k$.
Now, we establish that the setup we worked with earlier can be obtained as a (further) special case of the model in Section 7. First, we set $T_{i}=S\times\mathbb{E}_{i}$ so that the type of each agent is defined. $\hat{E}_{i}(s,E_{i})$ is therefore simply given by $E_{i}$. Note that $p$ was defined as a prior in the previous setting, whereas here $q_{i}$ represents a posterior belief once an agent is informed of her type. Since evidence is uncorrelated, we have
[…]
The map $\hat{\theta}_{i}$ now maps $S\times\mathbb{E}_{i}$ to $\mathbb{E}_{i}$ as we continue to assert constant preferences and with the aforementioned cost structure, only evidence is payoff relevant. The social choice function is defined as follows.
[…]
This definition does not lead to difficulties since only such profiles occur with positive probability where $s_{i}=s$ for all $i$. Therefore, we have derived the previous setup as a simplification of this one.
Now we will prove that evidence incentive compatibility is necessary for rationalizable implementation with arbitrarily small transfers. If $f$ is rationalizably implementable with arbitrarily small transfers, there is a mechanism $\mathcal{M}=(M,g,\tau)$ which implements $f$ such that for any profile of rationalizable messages $m\in R^{G}\left(t\right)$ of the game $G\left(\mathcal{M},\mathcal{T},u\right)$, we have $g(m)=f(t)$. Further, $\mathcal{M}$ can be chosen so that $|\tau_{i}(m)|\leq\frac{\varepsilon}{2}$ for every message profile $m$. The game $G\left(\mathcal{M},\mathcal{T},u\right)$ must also have a Bayesian Nash equilibrium. Such an equilibrium also must have outcome $f(t)$ (transfers notwithstanding), since every Nash Equilibrium survives the iterated elimination of strictly dominated strategies and thus constitutes a rationalizable message profile. Suppose this equilibrium is given by $\sigma$.  Consider an alternate mechanism $\mathcal{M}^{\prime}$ where agents merely report their type $t_{i}$ including the evidence and the designer plays their strategy from $\sigma$ on their behalf. The outcome and transfers are chosen to be identical to those induced by $\mathcal{M}$. Such a mechanism must have a truth-telling equilibrium, say $\sigma^{\prime}$. That is, for each agent $i$, and each type $t_{i}$, this type must have the incentive to truthfully report her type $t_{i}$ if she knows others are reporting their type accurately. That is,
$\sum_{t_{-i}\in T_{-i}}q_{i}(t_{i})(t_{-i})u_{i}(f(t_{i},t_{-i}),\tau_{i}(t_{i},t_{-i}),(t_{i},t_{-i}))\geq\sum_{t_{-i}\in T_{-i}}q_{i}(t_{i})(t_{-i})u_{i}(f(t_{i}^{\prime},t_{-i}),\tau_{i}(t_{i}^{\prime},t_{-i}),(t_{i},t_{-i}))\text{.}$
The above must be satisfied with for all $\varepsilon>0$. Then, taking limits with $\varepsilon\rightarrow 0$, we obtain
$\sum_{t_{-i}\in T_{-i}}q_{i}(t_{i})(t_{-i})\bar{u}_{i}(f(t_{i},t_{-i}),(t_{i},t_{-i}))\geq\sum_{t_{-i}\in T_{-i}}q_{i}(t_{i})(t_{-i})\bar{u}_{i}(f(t_{i}^{\prime},t_{-i}),(t_{i}^{\prime},t_{-i}))\text{,}$
which shows that $f$ satisfies evidence incentive compatibility.
The sufficiency proof is by construction of an implementing mechanism. Consider the following mechanism:
Message Space: $M_{i}=\mathbb{E}_{i}\times T_{i}\times T_{i}\times...\times T_{i}$ ($\bar{k}+J+1$) times, where $\bar{k}$ is the minimum order of belief at which any pair of states with different $f$-optimal outcomes can be separated by some agent. A typical message for agent $i$ is denoted by  $m_{i}=(E_{i},(t_{i}^{0,k})_{k=1}^{\bar{k}+1},(t_{i}^{j})_{j=1}^{J})$. Accordingly, $m=(m_{i})_{i\in\mathcal{I}}$ is a profile of messages.
Outcome: The outcome is given by $\frac{1}{J}\sum_{j=1}^{J}f(t^{j})$.
Transfers:
The mechanism uses the following transfers, which we document below. Let $\beta$ is a small positive number. Define
$\tau_{i}^{1}\left(m\right)=\beta\times|E_{i}|\text{.}$
This transfer incentivizes agents to submit all their evidence. Further, define
$\tau_{i}^{2,1}\left(m\right)=\beta\times[2\hat{\pi}_{i}^{1}(t_{i})(E_{-i})-\hat{\pi}_{i}^{1}(t_{i})\cdot\hat{\pi}_{i}^{1}(t_{i})]\text{;}$
$\tau_{i}^{2,k}\left(m\right)=\beta\times[2\hat{\pi}_{i}^{k}(t_{i}^{k})(\hat{\pi}_{-i}^{k-1}(t_{-i}^{k-1}),\hat{\pi}_{-i}^{k-2}(t_{-i}^{k-1}),...,\hat{\pi}_{-i}^{0}(t_{-i}^{k-1}))-\hat{\pi}_{i}^{k}(t_{i}^{k})\cdot\hat{\pi}_{i}^{k}(t_{i}^{k})],\forall k=2,...,\bar{k}+1.$
This transfer is a set of proper scoring rules which incentivizes agents to predict other agent’s lower order beliefs and their types using the report $t_{i}^{k}$. Given the above transfer definitions, $\beta$ can be chosen to be as small as required to meet any prescribed transfer bound $\varepsilon$. This is critical, since these transfers are active for any profile of rationalizable strategies.
From the above transfers, it is easy to see that the strategy that maximizes agent $i$’s utility from $\tau^{1}$ and $\tau^{2}$ is to present all her evidence and report truthfully in each $t_{i}^{0,k}$.
Given $\beta$, define as $\bar{\beta}_{i}$ the minimum loss in utility over any type profile that agent $i$ endowed with any evidence can induce owing to $\tau^{1}$ and $\tau^{2}$ by changing either her evidence report or her first $\bar{k}+1$ type reports to be different from the truth.
Now, we define the transfers used to discipline the last $J$ reports, which are used to set the outcome. The following transfer penalizes an agent for being among the first deviants from $t^{0,\bar{k}+1}$.
[…]
Finally, we define the following check between an agent’s report $t_{i}^{0,\bar{k}+1}$ and $t_{i}^{j}$.
[…]
where $\bar{\tau}^{4}>0$. This transfer is repeated for every round.
We scale the transfers as follows. First, notice that for any given $\varepsilon$, $\beta>0$ can be chosen so that the overall value of the transfers $\tau^{1}$ and $\tau^{2}$ does not exceed $\varepsilon$. This fixes each $\bar{\beta}_{i}$, and thereby $\min_{i}\bar{\beta}_{i}$. We now choose a large enough number of rounds ($J$) such that $\min_{i}\bar{\beta}_{i}>\frac{1}{J}$. This allows us to choose $\bar{\tau}^{3}$ such that $\min_{i}\bar{\beta}_{i}>\bar{\tau}^{3}>\frac{1}{J}$. Now, since $\bar{\tau}^{4}$ is only required to be positive, a sufficiently small value can be chosen so that $\min_{i}\bar{\beta}_{i}>\bar{\tau}^{3}+J\bar{\tau}^{4}>\frac{1}{J}$.
Now, we will prove that this mechanism implements $f$ in rationalizable strategies with arbitrarily small transfers. In the following proof, we denote the true state by $t^{\ast}=(t_{i}^{\ast})_{i\in\mathcal{I}}$.
Claim 12
Each agent presents all her evidence, i.e. $\hat{E}_{i}(t_{i}^{\ast})$.
Proof. If not, the agent can improve her payoff by providing additional articles of evidence due to the reward from $\tau^{1}$. There are no other incentives which are affected by the provision of additional evidence.
Claim 13
Each agent $i$ only presents such reports in $(t_{i}^{0,k})_{k=1}^{\bar{k}+1}$ which induce the same $k^{th}$ order beliefs as $t_{i}$ for any $k\leq\bar{k}+1$. Further, $f(t^{0,\bar{k}+1})=f(t)$.
Proof. From Claim 12, each agent presents all her evidence. Any report $t_{i}^{0,k}$ for $k<\bar{k}+1$, only controls the agent’s payoff from $\tau^{2}$ and therefore the only rationalizable report for any agent is to present a report $t_{i}^{0,k}$ which induces the same belief as the agent’s true type $t_{i}$. Any other report leads to a strictly lower payoff because $\tau^{2}$ is a quadratic scoring rule and yields a unique maximum.
Consider the agent’s report $t_{i}^{0,\bar{k}+1}$. It controls the agent’s payoff from $\tau^{2}$, $\tau^{3}$, and $\tau^{4}$ Changing $t_{i}^{0,\bar{k}+1}$ leads to a loss of at least $\min_{i}\bar{\beta}_{i}$, which by construction dominates gains from $\tau^{3}$ and $\tau^{4}$ (recall that $\min_{i}\bar{\beta}_{i}>J\bar{\tau}^{4}+\bar{\tau}^{3}$), so that it is not profitable to lie in $t_{i}^{0,\bar{k}+1}$. Higher-order measurability then yields that the reports in $t^{0,\bar{k}+1}$ yield the correct cell of the partition of $T$ to which $t$ belongs.
Claim 14
For any $j$, $t_{i}^{j}$ is the truth. Further, the mechanism implements and for any profile of rationalizable messages $m$, $\tau_{i}^{3}\left(m\right)=0$ and $\tau_{i}^{4}\left(m\right)=0$ for all $i$.
Proof. The proof is by induction. Consider any report $t_{i}^{j}$, $j\geq 1$ and suppose agents report truthfully all the way up to the ($j-1$)th message. If all agents $k\neq i$ report truthfully in $t^{j}$, then evidence incentive compatibility ensures that truth-telling is among the best responses for agent $i$. The small penalty from $\tau^{4}$ makes it the unique best response since from Claim 13, $t^{0,\bar{k}+1}$ is true. If other agents $k\neq i$ also lie in $t^{j}$, then the truth yields the agent a profit of $\bar{\tau}^{3}$ from $\tau^{3}$, which dominates any gains from affecting the outcome with a probability $\frac{1}{J}$, since $\bar{\tau}^{3}>\frac{1}{J}$. Therefore, such a message is not rationalizable. Therefore, it is rationalizable to present the truth in $t^{j}$. Then, the transfers $\tau^{3}$ and $\tau^{4}$ are clearly inactive. Further, if every $t^{j}$ is true, then the outcome is chosen correctly.